### v1

# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit anatomical and other individual differences, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results show that representations learned with these objectives generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data. In addition, we set new benchmarks for two foundational speech decoding tasks. Taken together, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data for speech decoding. We use abundant unlabelled neuroimaging data to pre-train a neural network using tasks that do not require explicit labels. The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation. We fine-tune the pre-trained network on labelled data, achieving faster generalisation owing to the strength of the representation learned in pre-training.
1 Introduction
In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases [50]. In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains [29, 7, 42, 46].
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present (e.g. [17]), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields (e.g. [27]), holds the promise of training models on collective, internet-scale data.
Given the scarcity of labelled data, self-supervised learning (SSL) appears promising as an avenue for domains where such data is rare or hard to obtain [3]. In the SSL paradigm, pretext tasks pre-train a model on unlabelled data by generating implicit training labels through transformations of the input data in order to help a downstream task. We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals. In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners. We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive magnetoencephalography (MEG), setting the baselines for speech detection and voicing classification on this data. The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time. Our main contributions are:
  • A set of domain specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
  • A neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
  • A comprehensive experimental evaluation providing evidence that our approach can scale up with data and enable cross-dataset, task, and subject generalisation.
2 Related Work
Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments. This is true both in recent state-of-the-art invasive studies [40, 39, 60, 9] and non-invasive studies [52]. These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments. Focusing on semantic rather than phonetic decoding, work by Tang et al. [52] is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos. They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions. Csaky et al. [11] find that a subset of individuals perform better when evaluated with a group-level model than with individual models. Exploiting audio data in a multi-modal framework, Défossez et al. [17] show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated. Although they repeat the result within two MEG and two EEG datasets, Défossez et al. [17] do not show any improvements for pooling data across datasets. Moreover, they do not combine data from studies with different labels either; cf. [58, 16, 56]. Unfortunately, two of these papers [58, 16] included a bug in their evaluation code. As such, their methods may perform no better than a baseline that provides pure noise inputs to the model [28].
[…]
In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning [1, 15, 41, 31, 62, 21]. Until now, similar approaches have not translated to the brain decoding literature. However, prior work has used other methods to leverage unlabelled brain data. For example, Jiang et al. [27] succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective. Although this work combined unlabelled EEG datasets, it only achieved improvements on non-speech tasks. Wang et al. [57] used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations. They attained impressive speech detection results but with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data.
3 Method
To process continuous neuroimaging data, we introduce a neural architecture for encoding heterogeneous brain signals into latent representations. By developing pretext tasks with the objective of learning generalisable brain representations, we leverage this architecture for self-supervised learning from unlabelled data, hoping to replicate similar successes in computer vision [21, 8].
3.1 Network Architecture
We design a two-stage neural network architecture (Figure 2). The pre-training stage uses pretext tasks to train a representation with unlabelled brain data. Then, the fine-tuning stage uses this representation to learn the downstream task by fine-tuning with labelled data.
Figure 2: Architecture overview. Inputs are projected into a shared dimension, then encoded. In pre-training, all weights are trainable except for modules in light-red, while in shallow fine-tuning, modules with light-blue borders are frozen and modules with light-red borders are unfrozen. Deep fine-tuning is identical, except the encoder is trainable. Dashed borders indicate optional components.
Normalisation. We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance.
Backbone. The network takes as input the standardised sample windows. To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$. Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs [61, 14]. Specifically, our convolutional encoder adapts the implementation of the SEANet architecture [51] used in Défossez et al. [14]. As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions. We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$. Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings (where $\tau<t$), each of dimension $d_{\mathrm{backbone}}$ (i.e. an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
Pre-training. Following the advice of Balestriero et al. [3, Section 3.2], we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation. After the projector, linear classifiers make predictions for each of the pretext tasks.
Fine-tuning. In this stage, we introduce classifiers for the downstream tasks and train with labelled data and supervised learning (Figure 2). Depending on the experiment, we fine-tune in one of two ways. In the first, we train an MLP classifier from scratch on top of the pre-trained representation, which remains frozen. Thus, we backpropagate only through the classifier. We call this shallow fine-tuning. In the other, we fine-tune by training an MLP on top of the pre-trained representation again, but do not freeze the backbone, training end-to-end. Here, we backpropagate through the classifier and the backbone. We refer to this as deep fine-tuning. In both cases, a new trainable dataset-specific linear layer can be introduced for a novel dataset.
For speech detection, our classifier makes a prediction for each individual embedding. For voicing classification, where there is only one label for each sample window, the embeddings are flattened into a tensor in $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ representing the entire window. This is the input to the voicing classifier and is referred to as full epoch decoding in neuroimaging literature [12].
Subject conditioning. Just as speakers have different voices, neural responses between subjects have different characteristics. Consequently, individual variation leads to models that do not generalise well across subjects [11]. In the speech literature, models include speaker conditioning to account for these differences [20]. We take a similar approach by introducing subject conditioning, represented as a subject embedding, to our model. With a SEANet-based architecture, Zeghidour et al. [61] find that conditioning is equally as effective at the encoder bottleneck as in other stages of the model. Hence, we place ours at the cortex encoder bottleneck for simplicity. We evaluate multiple types of conditioning including subject embeddings and feature-wise linear modulation (FiLM) [44].
3.2 Pretext Tasks
Our pretext tasks are unsupervised feature learning tasks for continuous neuroimaging signals that aim to learn generalisable speech decoding features. Since different datasets use different hardware and varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal. This means that these tasks do not require identifying specific sensor channels.
Band prediction. In the literature, neural responses can be broadly segmented into functional frequency bands [22, 45, 36]. Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech [35], Theta ($\theta$) waves (4–8 Hz) reliably track [34] and phase-lock to the amplitude envelope of heard sentences [43], Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals [49], Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding [5] which affects lexical processing [59], Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g. memory, learning, reasoning, and planning) [19, 6], and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection [26] and phonemic feature classification in the STG [38] as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) [10]. As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected. This yields the loss
$\mathcal{L}_{\mathrm{band}}=\sum_{x\in B}\mathcal{L}_{\mathrm{CE}}(f_{\mathrm{band}}(g(x^{\omega^{\prime}})),\omega),$ (1)
where $B$ is a mini-batch of samples, $\omega\in\{\delta,\theta,\alpha,\beta,\gamma,\gamma^{\mathrm{high}}_{\mathrm{lower}},\gamma^{\mathrm{high}}_{\mathrm{upper}}\}$, and $\mathcal{L}_{\mathrm{CE}}$ is the cross-entropy loss as this is a multi-class classification task.
Phase shift prediction. Phase coupling between networks of neuron populations is necessary for coordinating brain activity [18, 54]. Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature. Supporting this insight, recent work [27] also finds phase to be an essential component of the signal.
To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho$ of the sensors. Applying this shift to random sensors is critical since sensors are placed in different positions, capturing different regions of the brain. Uniform random selection ensures differences between any two regions of the brain are represented. The objective of this task is to predict the phase shift. This leads to a similar loss
$\mathcal{L}_{\mathrm{phase}}=\sum_{x\in B}\mathcal{L}_{\mathrm{CE}}(f_{\mathrm{phase}}(g(x^{\phi})),\phi),$ (2)
where $x^{\phi}$ describes the signal with a phase shift $\phi$ applied to a proportion of the sensors. We use a discrete number of possible phase shifts, treating it as a multi-class task rather than a regression task, to ease the difficulty of the problem as MEG scanners typically have a large number of sensors.
Amplitude scale prediction. MEG and EEG signals use an array of sensors at different spatial locations, capturing different signal sources more intensely. Representing the relative amplitude difference between sensors could be important for differentiating between neural responses originating from distinct parts of the brain. Within speech, Hamilton et al. [26] find that localised regions of the STG respond to sustained speech and speech onsets. Differentiating between neural responses from this region and others may be essential for decoding speech perception.
Thus, this pretext task focuses on learning representations that encode relative sensor amplitude differences. Similar to the phase shift task, we select a random proportion of the sensors $\rho$ and apply a discrete random amplitude scaling coefficient $A\in[-2,2]$, discretised into 16 scaling factors, to the signal. The objective is to predict the scaling factor, leading to the loss
$\mathcal{L}_{\mathrm{amplitude}}=\sum_{x\in B}\mathcal{L}_{\mathrm{CE}}(f_{\mathrm{amplitude}}(g(x^{A})),A),$ (3)
where $x^{A}$ is the signal scaled with $A$.
These pretext tasks capture complementary time- and frequency-domain properties of the signal. Hence, during pre-training, we combine them, creating an augmented version of the input for every pretext task by applying the matching transformation. We feed the augmented inputs through the network backbone and apply the corresponding classifier to predict the transformation, summing the weighted losses such that our final pre-training loss is given by
$\mathcal{L}_{\mathrm{SSL}}=w_{1}\mathcal{L}_{\mathrm{band}}+w_{2}\mathcal{L}_{\mathrm{phase}}+w_{3}\mathcal{L}_{\mathrm{amplitude}},$ (4)
where $w_{i}$ is a constant coefficient for each self-supervised loss.
4 Experiments
In this section, we evaluate the representations learned with our pretext tasks by measuring their ability to scale downstream performance with unlabelled data. This includes understanding how well they can generalise across datasets, subjects, and tasks. We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG [32] and faster sampling rates than fMRI [25].
4.1 Experimental setup
Datasets. Unless specified otherwise, our experiments use Cam-CAN [48, 53] as an unlabelled representation learning dataset for pre-training. This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings. For our downstream tasks, we use two labelled heard speech MEG datasets where participants listen to short stories or audiobooks. Armeni et al. [2] contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al. [24] has 27 subjects, each recorded for 2 hours (54 hours total). Overall, we utilise over 200 hours of data. To the best of our knowledge, this is the largest volume of MEG data ever used for speech decoding.
Preprocessing. Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner. To eliminate high-frequency muscle movement artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.5Hz to remove slow-drift artifacts. Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise. Next, we downsample the signal to 250Hz, avoiding aliasing at frequencies up to our low-pass filter threshold. Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
Downstream tasks. We evaluate our methods with two fundamental speech decoding tasks of increasing difficulty. The first, speech detection, determines whether speech occurs in the auditory stimulus using the neural response. The second task is voicing classification. Given data aligned at the occurrence of a phoneme, the task is to recognise whether the phoneme is voiced or voiceless, where voicing is a binary phonetic feature that categorises whether a speech sound is associated with vocal cord vibration. We select these tasks as they are simpler than phoneme recognition, but are foundational because they must be solved to decode speech accurately into natural language.
Training. We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation. Appendix B provides complete training details for all experiments.
4.2 Learning Generalisable Representations Using Pretext Tasks
To understand whether the pretext tasks produce useful representations, we investigate how they generalise independently with shallow fine-tuning (i.e. with the representation frozen). In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, while for voicing the band prediction task does best. In all cases, the tasks either match or beat a randomly initialised and shallow fine-tuned baseline. The results are statistically significant whenever we beat the baselines and provide initial evidence that all of our tasks are useful for learning representations that generalise downstream to speech decoding.
Pretext task Fine-tuned balanced accuracy
Speech $t$ $p$ Voicing $t$ $p$
None (random backbone) $0.500\pm 0.000$- -- $0.500\pm 0.000$- --
Phase($\rho=0.5$) $0.538\pm 0.002$ $18.5$ $0.001$ $0.500\pm 0.000$- --
Amplitude($\rho=0.2$) $\mathbf{0.555}\pm 0.002$ $23.3$ $0.001$ $0.500\pm 0.000$- --
Band $0.551\pm 0.005$ $10.0$ $0.005$ $\mathbf{0.503}\pm 0.001$ $4.1$ $0.03$
Chance-level $0.5$- -- $0.5$- --
Table 1: Pre-training with pretext tasks produces generalisable representations. We provide the test accuracy at the point of best validation accuracy (early stopping). In the none baseline, the network backbone is randomly initialised before fine-tuning (i.e. there is no pre-training). The uncertainty is the standard error of the mean over three random seeds. We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy. Models are trained on all subjects in Armeni et al. [2] and evaluated on held-out sessions.
The strength of band prediction in voicing classification may be due to discriminating between theta waves and other frequency bands as theta waves are associated with phonetic feature processing. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix C). We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset [26]. In contrast, a larger $\rho$ for phase shift prediction may be best because phase is descriptive of neural synchrony which is distributed across the brain. Unlike localised responses, a large proportion of the sensors can detect this feature.
4.3 Scaling Speech Decoding With Unlabelled Data
Previously, we pre-trained and fine-tuned using data from Armeni et al. [2]. This setting was useful to understand whether our pretext tasks benefited the representation but did not demonstrate if they were effective in a real scarce data setting, where representations must generalise across datasets in order to scale. Furthermore, we did not investigate the scaling behaviour with downstream tasks. Here, we pre-train with Cam-CAN [48, 53], which does not have any speech decoding labels, and fine-tune for ten epochs on Armeni et al. [2] and Gwilliams et al. [24], showcasing transfer with minimal fine-tuning. This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI. We use shallow fine-tuning for speech detection and deep fine-tuning for voicing. As evidenced by Table 1, we require the latter because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
Figure 3: Scaling unlabelled data improves downstream generalisation. We pre-train the model on increasing amounts of unlabelled data and fine-tune on voicing and speech detection. No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area shows the standard error of the mean across three seeds.
Figure 3 shows balanced accuracy as we increase the amount of unlabelled data in pre-training up to approximately 160 hours. For both tasks, pre-training outperforms the no pre-training baseline, which fails to generalise at all, showing that any unlabelled data is sufficient to pre-train a useful representation. In both tasks, there is a clear improvement in accuracy as the amount of unlabelled data increases. Thus, adding unlabelled data improves generalisation and scales performance. With speech detection, the improvement in accuracy is more significant for Armeni et al. [2]. Noting that there is more within-subject data in this dataset than Gwilliams et al. [24], speech detection may be more sensitive to subject variation than voicing classification.
We scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our self-supervised method is an exception to the common consensus that pooling subjects worsens generalisation. Moreover, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly notable as the Armeni et al. [2], Gwilliams et al. [24], and pre-training datasets all use different scanners entirely. Performing well across these datasets indicates that our pretext tasks successfully generate representations that are generalisable across heterogeneous scanners. Beyond hardware, our results also show improvements on distinct tasks, thus achieving cross-task generalisation. Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks, showing that remarkably any unlabelled data can be used to scale.
Whilst we have shown that our representations generalise faster than without pre-training, do they outperform supervised baselines when trained to saturation? While voicing requires deep fine-tuning to see a substantial learning signal, the speech detection task, with which we use shallow fine-tuning, can be compared directly to similarly parameterised supervised classifiers. Table 2 provides this comparison. A representation pre-trained with 160 hours of unlabelled data significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets. Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is much larger without an encoder. Even with this bias towards the supervised classifier, the self-supervised case performs better. Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
Model Speech detection balanced accuracy
Armeni $t$ $p$ Gwilliams $t$ $p$
Supervised
Linear $0.5222\pm 0.0015$ $12$ $0.003$ $0.5006\pm 0.0001$ $5$ $0.02$
Self-supervised + fine-tuning
Random backbone
   MLP (two layers) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
Pre-trained backbone
   Linear $0.5853\pm 0.0017$ $52$ $0.0003$ $0.5011\pm 0.0001$ $9$ $0.006$
   MLP (two layers) $\mathbf{0.6101}\pm 0.0017$ $51$ $0.0002$ $\mathbf{0.5230}\pm 0.0002$ $113$ $0.00004$
Chance-level $0.5$- -- $0.5$- --
Table 2: Self-supervised pre-training outperforms supervised classifiers. We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion. We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds. We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance.
4.4 Generic Self-Supervised Representations Improve Speech Decoding
Figure 4: Cross-task semi-supervised learning does not improve generalisation. We pre-train with increasing amounts of labelled data from Armeni et al. [2], using speech labels when evaluating voicing and voicing labels for speech. We also use a constant 38 hours of unlabelled data in pre-training. No pre-training shows test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). SSL is when pre-training with only the unlabelled data. The shaded area shows the standard error of the mean across three seeds.
Our self-supervised tasks led to clearly generalisable and scalable representations by providing an inductive bias towards speech decoding. However, could these representations be made more powerful by utilising labels from other speech decoding tasks to improve this bias? This moves us into the semi-supervised learning regime, where we use a mix of labelled and unlabelled data.
With the same experimental setup as Figure 3, we add speech detection or voicing as an additional loss during pre-training. Figure 4 shows that no amount of additional labelled data improves downstream accuracy, with all data points remaining near the range of the standard error of the mean of the SSL baseline for both tasks. There also does not appear to be any indication of scaling within the range of labelled data available to us.
This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy. The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space. Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks. This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech. However, the overall point remains salient. The generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
4.5 Generalising Zero-Shot To Novel Subjects
Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11]. Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge. We pre-train then fine-tune our models on a subset of Gwilliams et al. [24], holding out three subjects for evaluation.
We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation. We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation. For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
Subject Subject voicing balanced accuracy
conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
FiLM [44] $\mathbf{0.5197}\pm 0.001$ $18.6$ $0.001$ $\mathbf{0.5058}\pm 0.002$ $3.3$ $0.04$
Chance-level $0.5$- -- $0.5$- --
Table 3: Self-supervised representations generalise zero-shot to novel subjects. We quote test accuracy at the point of best validation accuracy (early stopping). Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
Table 3 reveals that in-distribution subject generalisation is helped by using subject conditioning methods. Both subject embeddings and FiLM conditioning outperform no conditioning, with FiLM being best. However, for novel subject generalisation, using a subject conditioning method makes little difference. Regardless of conditioning, we can generalise beyond chance-level, but suffer a reduction in accuracy compared to in-distribution subjects. Conditioning may be negligible in the unseen case as the embeddings are random rather than learned for the held-out subjects.
4.6 Limitations
Although our results are significant in demonstrating a viable path forward to scale up speech BCIs, there remain a number of limitations to the present work. We focused here on two downstream tasks: speech detection and voice classification. Ultimately, we would like to expand this work to predict full transcripts from brain recordings (i.e. brain-to-text). This has been achieved with surgical data [40, 60] but not yet convincingly with non-invasive methods like MEG or EEG [28]. Speech detection has played an important role in the development of full brain-to-text in a surgical context [40] and we hope may play a similar role for non-invasive methods. Prior work has further used voice classification as a stand in for phoneme classification [23], and we have been able to improve on these results here. In future work, we would like to expand this to all English phonemes. Secondly, while we have been able to demonstrate the utility of a few pretext tasks here, we do not claim to have exhausted the full set of useful tasks. Rather, we conjecture that more useful pretext tasks remain to be found and believe a useful avenue of research will be into other input representations for brain recordings. For example, this paper did not make use of spatial features. Another limitation is our emphasis on heard speech over other types of speech, such as attempted or imagined speech. We hypothesise that the same methods presented here will generalise to these other varieties of speech, though this has yet to be shown. But, perhaps the biggest limitation of the present work is that, while it surpasses the amount of data used in other studies, it remains to be seen how much speech decoding tasks can be improved by scaling up the number of datasets used in training. In sharing this work now, we believe that the current proof of concept will be sufficiently impactful to the field as we continue to actively scale up the datasets that we can leverage.
5 Conclusion
Ultimately, solving speech decoding could transform the lives of patients with severe communication difficulties. Yet, this promise has not materialised because the field has been blocked by its inability to scale up data to leverage deep learning. Prior methods have been unable to aggregate data across different datasets, labels, or subjects to scale up because of heterogeneity in recording hardware, experiment design, and participants. A handful of studies have shown weak signals towards alleviating these issues. But until now, no one has developed a general solution. We provided a unified method that leverages unlabelled data using generic pretext tasks that shows that all of these problems can be solved. We verified this with experiments showing that our method not only scales with heterogeneous data but even generalises across datasets, subjects, and tasks. Our method unlocks the potential of the bitter lesson, providing a general method to exploit more computation by using more data. We implore the research community to employ the vast quantities of data and compute available to realise this potential. If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
Acknowledgments and Disclosure of Funding
We would like to thank Botos Csaba for many insightful discussions and creative ideas which helped shaped the direction of this work. Thanks also to Minqi Jiang for an encouraging conversation on unsupervised representation learning, Mats W.J. van Es for technical assistance with the OSL library, Brian Liu for technical contributions which did not reach the final paper, and Miran Özdogan for reviewing a draft of this work. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558.
[…]
Appendix A Supervised Heard Speech Decoding
While this work is primarily about scaling speech decoding, we also set the first baselines for speech detection and voicing classification from MEG with connected auditory speech stimuli. Therefore, in this section we provide fully supervised results where our architecture is trained to completion using all available labelled data in the corresponding dataset (except the validation and test splits) and compare this to a linear classifier baseline. These results are not comparable with our self-supervised models because the entire architecture is trained end-to-end. In our SSL work with shallow fine-tuning, only the final classifier is trained with labels.
Experiment Balanced accuracy
Speech detection Voicing
Armeni (linear) $0.5222\pm 0.0015$ $0.5207\pm 0.0007$
Armeni (ours) $\mathbf{0.7253}\pm 0.0005$ $\mathbf{0.5282}\pm 0.0006$
Gwilliams (linear) $0.5006\pm 0.0001$ $0.5166\pm 0.0002$
Gwilliams (ours) $\mathbf{0.5946}\pm 0.0016$ $\mathbf{0.5211}\pm 0.0005$
Table 4: Supervised baselines. Uncertainty is standard error of the mean over three random seeds.
The results in Table 4 show that we can achieve strong speech detection results on Armeni et al. [2] and Gwilliams et al. [24] with our architecture. The difference between the two may be due to a combination of different scanners, numbers of sensors (208 vs 269 magnetometers), varied difficulty in the audio stimulus, and more within-subject data in Armeni et al. [2]. We also achieve good results on voicing classification for both datasets, easily outperforming the linear models.
One potential way to improve on both supervised and self-supervised approaches is by using a temporally aware model such as an LSTM or an attentional model such as a Transformer as the classifier. We do not investigate this within this paper as it is out of scope. Rather, we choose to use simple classifiers as the purpose of this paper is to demonstrate the potential of scaling.
Appendix B Experiment Details
To generate the data in Table 1, we pre-train our representations with non-overlapping sample windows from all subjects and sessions 001 to 009 of Armeni et al. [2] for 200 epochs with a $0.8:0.1:0.1$ train/val/test split. We shallow fine-tune with labels for 30 epochs using all subjects, but only session 010, which was held-out during pre-training.
For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting. We pre-train only with Shafto et al. [48], Taylor et al. [53]. We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include. Each seed uses a different set of subjects to reduce negative effects from outlier subjects. We fine-tune with all data in Armeni et al. [2] and Gwilliams et al. [24] with the same train/val/test ratios as before.
For the results shown in Figure 4, we use the same setup as in Figure 3. In addition to the unlabelled data, we add labelled data from Armeni et al. [2], increasing the number of sessions used in the sequence 1, 3, 6, 9, 12, and 27. For this experiment, we calculate the number of labelled data hours using an estimate of 1 hour per session per subject.
Finally, for the results in Table 3, we pre-train and deep fine-tune to saturation with subjects 01-07 of Gwilliams et al. [24] and test zero-shot generalisation using held-out subjects 08-10. All other experimental details are the same as the previous two experiments.
In all experiments, we use three randomly selected seeds for each pre-training and corresponding fine-tuning run. For speech detection, since our encoder reduces the temporal dimension from 125 samples (the number of samples in a 0.5 second window with a sample rate of 250Hz) down to 5 embeddings, we downsample our speech detection labels to match using PyTorch’s torch.nn.functional.interpolate. Therefore, each speech detection label represents a 0.1 second period of time.
Appendix C Hyperparameters
We conduct a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture. In all experiments in this section, we measure downstream performance (using shallow fine-tuning) with best validation accuracy for speech detection unless explicitly specified, training and evaluating only on data from subject 001 of Armeni et al. [2] to minimise training time, over three random seeds. All uncertainties in this section are standard deviations and not standard error of the mean as in the main body.
Table 5 shows the results of a search over the proportion $\rho$ of sensors to apply the phase and amplitude pretext task transformations to. We only search up to $\rho=0.5$ as anything above this is equivalent to applying the inverse transformation to a proportion $1-\rho$ of the sensors. We find that for phase shift prediction, $\rho=0.5$ is optimal, while for amplitude scale prediction $\rho=0.2$ is best.
Task $\rho$ Balanced accuracy
Phase shift prediction 0.1 $0.5452\pm 0.0048$
0.2 $0.5566\pm 0.0014$
0.3 $0.5578\pm 0.0021$
0.4 $0.5530\pm 0.0014$
0.5 $\mathbf{0.5655}\pm 0.0026$
Amplitude scale prediction 0.1 $0.5440\pm 0.0012$
0.2 $\mathbf{0.5796}\pm 0.0050$
0.3 $0.5625\pm 0.0063$
0.4 $0.5684\pm 0.0016$
0.5 $0.5650\pm 0.0026$
Table 5: Best $p$ for phase and amplitude pretext tasks.
In our architecture, we define $d_{\mathrm{backbone}}$ to be the bottleneck dimension of the cortex encoder (i.e. the embedding dimension) as well as the size of $d_{\mathrm{shared}}$. Balestriero et al. [3], using the results from work by Bordes et al. [4], suggest that increasing the backbone output dimension in self-supervised learning leads to better downstream transfer. We investigate this for our architecture by varying the encoder dimension in Table 6. Our results are consistent with their findings. Balanced accuracy is highest at the largest encoder dimension we evaluated. Encoder dimensions larger than $2048$ led to unstable training and were not feasible given the compute resources available to us.
$d_{\mathrm{backbone}}$ Balanced accuracy
$128$ $0.5775\pm 0.0037$
$300$ $0.5715\pm 0.0012$
$512$ $0.5835\pm 0.0016$
$1024$ $0.5800\pm 0.0050$
$2048$ $\mathbf{0.5987}\pm 0.0024$
Table 6: Optimal encoder dimension.
In Table 7, we show the effect of varying the length $w$, i.e. the window size of each data sample. Increasing the window to five seconds led to unstable losses in pre-training so we did not evaluate above a window length of two seconds. This is likely to be because increasing the window size reduces the number of training samples when only non-overlapping samples are used. For speech detection, there appears to be a U-shaped distribution, where very short and very long windows lead to better downstream accuracy. For voicing classification, the window lengths we evaluated seem to make little difference. However, windows smaller than 0.1 seconds may lead to significantly lower accuracy as most phonemes will occur for at least this length of time in the neural response [23].
Window length (s) Balanced accuracy
Speech detection Voicing classification
$0.2$ $0.5919\pm 0.0014$ $0.5033\pm 0.0014$
$0.4$ $0.5804\pm 0.0036$ $\mathbf{0.5077}\pm 0.0027$
$0.5$ $0.5715\pm 0.0012$ $0.5055\pm 0.0023$
$1.0$ $0.5790\pm 0.0012$ $0.5062\pm 0.0020$
$2.0$ $\mathbf{0.6116}\pm 0.0019$ $0.5065\pm 0.0011$
Table 7: Ideal sample window length.
We also evaluated several different configurations of the SEANet architecture for the cortex encoder, converging on convolutional blocks with channel dimensions $(512,512,512,512)$.
While these ablations indicate a theoretically ideal architectural configuration, in practice, we altered our final experimental architecture due to instabilities during training when data was scaled up. Our final architecture hyperparameters achieve a balance between the best values from our hyperparameter search and stable training. These values are detailed in Table 8.
Hyperparameter Value
Window length (s) $0.5$
$\rho$ (phase) $0.5$
$\rho$ (amplitude) $0.2$
$\{w_{1},w_{2},w_{3}\}$$\{1.0,1.0,1.0\}$
$d_{\mathrm{shared}}$ $512$
$d_{\mathrm{backbone}}$ $512$
SEANet convolution channels $(512,512,512,512)$
SEANet downsampling ratios $(5,5,1)$
FiLM conditioning dimension $16$
Subject embedding dimension $16$
Pre-training epochs $200$
Optimizer AdamW [33]
Learning rate $0.000066$
Train ratio $0.8$
Validation ratio $0.1$
Test ratio $0.1$
Table 8: Experimental hyperparameters.
Appendix D Compute Resources
All experiments were run on individual NVIDIA V100 GPUs with up to 32GiB of GPU memory on a system with up to 128GiB of RAM. Each pre-training run with the maximum amount of pre-training data took approximately 150 hours (6.25 days). Fine-tuning following pre-training took up to another 12 hours. We estimate that we used approximately 2000 hours of compute for the final experimental runs, including hyperparameter searches. In total, over the course of developing this work from idea to final paper, we used around 10,000 hours of GPU compute.
Appendix E Licences For Datasets And Code
The Armeni et al. [2] dataset is distributed under CC-BY-4.0 while the Gwilliams et al. [24] dataset is distributed under the CC0 1.0 Universal licence. The Schoffelen et al. [47] dataset is distributed with a RU-DI-HD-1.0 licence from the Donders institute. The licence for the Cam-CAN [48, 53] dataset is unknown. The SEANet code adapted from Défossez et al. [14] is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.
Appendix F Broader Impacts
Decoding speech from non-invasive brain recordings is likely to bring about significant positive societal impacts. Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g. stammering). As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants. Nevertheless, the maturity of this technology could also present potential negative societal impacts. For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals. Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent. Moreover, there are possible long horizon effects associated with speech decoding research. Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society. In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality. Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse. Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.

### v2

@@
 To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings.
-Experimental results show that representations learned with these objectives generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data.
+Experimental results show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data.
 In addition, we set new benchmarks for two foundational speech decoding tasks.
@@
 Figure 1: Leveraging unlabelled data for speech decoding.
-We use abundant unlabelled neuroimaging data to pre-train a neural network using tasks that do not require explicit labels.
+We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled neuroimaging data, permitting learning from large heterogeneous datasets.
 The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation.
@@
 In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains [29, 7, 42, 46].
-In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet fully assimilated.
+In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated.
 State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech.
@@
 We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation.
-Appendix B provides complete training details for all experiments.
+Appendix E provides complete training details for all experiments.
 4.2 Learning Generalisable Representations Using Pretext Tasks
-To understand whether the pretext tasks produce useful representations, we investigate how they generalise independently with shallow fine-tuning (i.e.
+We investigate whether our self-supervised objectives produce useful representations by first learning a representation with each pretext task independently and then analysing its downstream generalisation using shallow fine-tuning (i.e.
 with the representation frozen).
-In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, while for voicing the band prediction task does best.
-In all cases, the tasks either match or beat a randomly initialised and shallow fine-tuned baseline.
-The results are statistically significant whenever we beat the baselines and provide initial evidence that all of our tasks are useful for learning representations that generalise downstream to speech decoding.
-Pretext task Fine-tuned balanced accuracy
-Speech $t$ $p$ Voicing $t$ $p$
-None (random backbone) $0.500\pm 0.000$- -- $0.500\pm 0.000$- --
-Phase($\rho=0.5$) $0.538\pm 0.002$ $18.5$ $0.001$ $0.500\pm 0.000$- --
-Amplitude($\rho=0.2$) $\mathbf{0.555}\pm 0.002$ $23.3$ $0.001$ $0.500\pm 0.000$- --
-Band $0.551\pm 0.005$ $10.0$ $0.005$ $\mathbf{0.503}\pm 0.001$ $4.1$ $0.03$
+In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, and critically, outperform a randomly initialised and shallow fine-tuned baseline.
+The results are statistically significant in all cases, providing initial evidence that all of our tasks are helpful in learning representations that generalise downstream to speech decoding.
+Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own.
+As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding.
+For our other downstream task, voicing classification, we were not able to use shallow fine-tuning to generalise to a statistically significant degree with any of the pretext tasks nor with a random backbone.
+It may be too difficult to learn with shallow fine-tuning and so we turn to using deep fine-tuning for this task in the next section.
+Pretext task Speech detection balanced accuracy
+Armeni $t$ $p$ Gwilliams $t$ $p$
+Amp($\rho=0.2$) $0.5846\pm 0.0032$ $26$ $7\mathrm{e}{-4}$ $0.5116\pm 0.0004$ $29$ $6\mathrm{e}{-4}$
+Phase($\rho=0.5$) $0.5831\pm 0.0029$ $29$ $6\mathrm{e}{-4}$ $0.5093\pm 0.0014$ $7$ $1\mathrm{e}{-2}$
+Band $0.5941\pm 0.0024$ $39$ $2\mathrm{e}{-3}$ $0.5177\pm 0.0028$ $6$ $1\mathrm{e}{-2}$
+All $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
+None (rand. backbone) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
 Chance-level $0.5$- -- $0.5$- --
@@
 there is no pre-training).
+When all pretext tasks are used, their losses are weighted equally.
 The uncertainty is the standard error of the mean over three random seeds.
 We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
-Models are trained on all subjects in Armeni et al.
-[2] and evaluated on held-out sessions.
-The strength of band prediction in voicing classification may be due to discriminating between theta waves and other frequency bands as theta waves are associated with phonetic feature processing.
-The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix C).
+Models are pre-trained on all of Cam-CAN [48, 53] and fine-tuned on Armeni et al.
+[2] or Gwilliams et al.
+[24] respectively.
+Among the pretext tasks, band prediction performs best.
+This could be because the representation easily identifies phase-locking to speech onset in theta waves [43].
+However, further investigation is necessary to determine which frequency band is most informative.
+The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix F).
 We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors.
@@
 Unlike localised responses, a large proportion of the sensors can detect this feature.
+Whilst we have shown that our representations generalise, do they outperform supervised baselines when trained to saturation?
+Using shallow fine-tuning once more, we compare our approach directly to similarly parameterised supervised classifiers in Table 2.
+A representation pre-trained with all pretext tasks significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets.
+Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is far larger without an encoder.
+Even with this bias favouring the supervised classifier, the self-supervised case performs better.
+Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
+Model Speech detection balanced accuracy
+Armeni $t$ $p$ Gwilliams $t$ $p$
+Supervised
+Linear $0.5222\pm 0.0015$ $12$ $3\mathrm{e}{-3}$ $0.5006\pm 0.0001$ $5$ $2\mathrm{e}{-2}$
+Self-supervised + fine-tuning
+Random backbone
+   MLP (two layers) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
+Pre-trained backbone
+   Linear $0.5853\pm 0.0017$ $52$ $3\mathrm{e}{-4}$ $0.5011\pm 0.0001$ $9$ $6\mathrm{e}{-3}$
+   MLP (two layers) $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
+Chance-level $0.5$- -- $0.5$- --
+Table 2: Self-supervised pre-training outperforms supervised classifiers.
+We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion.
+We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds.
+We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance.
+Self-supervised models are pre-trained on all of Cam-CAN [48, 53] and shallow fine-tuned on Armeni et al.
+[2] or Gwilliams et al.
+[24] while the supervised models are trained entirely on the latter datasets.
 4.3 Scaling Speech Decoding With Unlabelled Data
-Previously, we pre-trained and fine-tuned using data from Armeni et al.
-[2].
-This setting was useful to understand whether our pretext tasks benefited the representation but did not demonstrate if they were effective in a real scarce data setting, where representations must generalise across datasets in order to scale.
-Furthermore, we did not investigate the scaling behaviour with downstream tasks.
-Here, we pre-train with Cam-CAN [48, 53], which does not have any speech decoding labels, and fine-tune for ten epochs on Armeni et al.
+Figure 3: Scaling unlabelled data improves downstream generalisation.
+We pre-train the model on increasing amounts of unlabelled data from Cam-CAN [48, 53] and fine-tune on voicing and speech detection on the Armeni et al.
 [2] and Gwilliams et al.
-[24], showcasing transfer with minimal fine-tuning.
-This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI.
-We use shallow fine-tuning for speech detection and deep fine-tuning for voicing.
-As evidenced by Table 1, we require the latter because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
-Figure 3: Scaling unlabelled data improves downstream generalisation.
-We pre-train the model on increasing amounts of unlabelled data and fine-tune on voicing and speech detection.
+[24] datasets.
 No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
@@
 The shaded area shows the standard error of the mean across three seeds.
+Previously, we showed that the combination of all pretext tasks leads to the best representations and that this outperforms a randomly initialised backbone as well as a supervised classifier.
+In this section, we examine how these pretext tasks fare as we increase the amount of unlabelled data, analysing scaling performance on downstream tasks.
+As before, we pre-train with the combined pretext tasks using Cam-CAN [48, 53], which does not have any speech decoding labels, and now fine-tune for ten epochs on Armeni et al.
+[2] and Gwilliams et al.
+[24], showcasing transfer with minimal fine-tuning.
+This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI.
+Where earlier we analysed only speech detection, here, we also focus on voicing classification by using deep fine-tuning for this task only.
+As evidenced by Section 4.2, we require this because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
 Figure 3 shows balanced accuracy as we increase the amount of unlabelled data in pre-training up to approximately 160 hours.
@@
 [24], and pre-training datasets all use different scanners entirely.
-Performing well across these datasets indicates that our pretext tasks successfully generate representations that are generalisable across heterogeneous scanners.
-Beyond hardware, our results also show improvements on distinct tasks, thus achieving cross-task generalisation.
-Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks, showing that remarkably any unlabelled data can be used to scale.
-Whilst we have shown that our representations generalise faster than without pre-training, do they outperform supervised baselines when trained to saturation?
-While voicing requires deep fine-tuning to see a substantial learning signal, the speech detection task, with which we use shallow fine-tuning, can be compared directly to similarly parameterised supervised classifiers.
-Table 2 provides this comparison.
-A representation pre-trained with 160 hours of unlabelled data significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets.
-Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is much larger without an encoder.
-Even with this bias towards the supervised classifier, the self-supervised case performs better.
-Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
-Model Speech detection balanced accuracy
-Armeni $t$ $p$ Gwilliams $t$ $p$
-Supervised
-Linear $0.5222\pm 0.0015$ $12$ $0.003$ $0.5006\pm 0.0001$ $5$ $0.02$
-Self-supervised + fine-tuning
-Random backbone
-   MLP (two layers) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
-Pre-trained backbone
-   Linear $0.5853\pm 0.0017$ $52$ $0.0003$ $0.5011\pm 0.0001$ $9$ $0.006$
-   MLP (two layers) $\mathbf{0.6101}\pm 0.0017$ $51$ $0.0002$ $\mathbf{0.5230}\pm 0.0002$ $113$ $0.00004$
-Chance-level $0.5$- -- $0.5$- --
-Table 2: Self-supervised pre-training outperforms supervised classifiers.
-We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion.
-We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds.
-We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance.
-4.4 Generic Self-Supervised Representations Improve Speech Decoding
-Figure 4: Cross-task semi-supervised learning does not improve generalisation.
-We pre-train with increasing amounts of labelled data from Armeni et al.
-[2], using speech labels when evaluating voicing and voicing labels for speech.
-We also use a constant 38 hours of unlabelled data in pre-training.
-No pre-training shows test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
-without pre-training).
-SSL is when pre-training with only the unlabelled data.
-The shaded area shows the standard error of the mean across three seeds.
-Our self-supervised tasks led to clearly generalisable and scalable representations by providing an inductive bias towards speech decoding.
-However, could these representations be made more powerful by utilising labels from other speech decoding tasks to improve this bias?
-This moves us into the semi-supervised learning regime, where we use a mix of labelled and unlabelled data.
-With the same experimental setup as Figure 3, we add speech detection or voicing as an additional loss during pre-training.
-Figure 4 shows that no amount of additional labelled data improves downstream accuracy, with all data points remaining near the range of the standard error of the mean of the SSL baseline for both tasks.
-There also does not appear to be any indication of scaling within the range of labelled data available to us.
-This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy.
-The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space.
-Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks.
-This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech.
-However, the overall point remains salient.
-The generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
-4.5 Generalising Zero-Shot To Novel Subjects
+Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners.
+Beyond hardware, our results also show improvements on distinct tasks, namely speech detection and voicing classification, thus illustrating cross-task generalisation.
+Thus, our pretext tasks are sufficiently generic to produce representations that generalise to multiple downstream tasks, indicating their applicability to speech decoding more generally.
+Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks.
+Remarkably, this shows that brain data collected from any task (including those that are not linguistic) can be used to scale downstream speech decoding performance.
+4.4 Scaling Unlabelled Data Improves Zero-Shot Generalisation To Novel Subjects
 Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11].
 Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge.
-We pre-train then fine-tune our models on a subset of Gwilliams et al.
-[24], holding out three subjects for evaluation.
-We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation.
-We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation.
-For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
-Subject Subject voicing balanced accuracy
-conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
-None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
-Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
-FiLM [44] $\mathbf{0.5197}\pm 0.001$ $18.6$ $0.001$ $\mathbf{0.5058}\pm 0.002$ $3.3$ $0.04$
-Chance-level $0.5$- -- $0.5$- --
-Table 3: Self-supervised representations generalise zero-shot to novel subjects.
-We quote test accuracy at the point of best validation accuracy (early stopping).
-Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
-Table 3 reveals that in-distribution subject generalisation is helped by using subject conditioning methods.
-Both subject embeddings and FiLM conditioning outperform no conditioning, with FiLM being best.
-However, for novel subject generalisation, using a subject conditioning method makes little difference.
-Regardless of conditioning, we can generalise beyond chance-level, but suffer a reduction in accuracy compared to in-distribution subjects.
-Conditioning may be negligible in the unseen case as the embeddings are random rather than learned for the held-out subjects.
+This is critical in order to widely deploy speech BCIs for new patients.
+In the following experiment, we pre-train using the same setup as Figure 3, but fine-tune only on Gwilliams et al.
+[24].
+We use FiLM conditioning as this performs best for seen and unseen subjects (Appendix B).
+During training, we hold out three subjects with which we evaluate novel subject generalisation (i.e.
+unseen subject performance).
+Figure 4: Scaling pre-training data improves novel subject generalisation.
+As before, we pre-train with increasing amounts of unlabelled data from Cam-CAN [48, 53], but now fine-tune only on Gwilliams et al.
+[24].
+In the seen case, we evaluate on held-out sessions from subjects that were used in the training data.
+For the unseen case we evaluate on three completely held-out subjects.
+No pre-training shows the best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
+without pre-training).
+We show only one no pre-training line as the result is the same for both datasets.
+The shaded area is the standard error of the mean over three seeds.
+Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive trend in performance for unseen (novel) subjects.
+This is a promising result, indicating that scaling with our method is an encouraging direction for resolving the issues related to individual variance faced by prior work.
 4.5 Limitations
@@
 Rather, we choose to use simple classifiers as the purpose of this paper is to demonstrate the potential of scaling.
-Appendix B Experiment Details
+Appendix B Subject Conditioning
+To investigate subject conditioning approaches, we pre-train then fine-tune our models on a subset of Gwilliams et al.
+[24], holding out three subjects for evaluation.
+We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation.
+We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation.
+For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
+Subject Subject voicing balanced accuracy
+conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
+None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
+Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
+FiLM [44] $\mathbf{0.5197}\pm 0.001$ $18.6$ $0.001$ $\mathbf{0.5058}\pm 0.002$ $3.3$ $0.04$
+Chance-level $0.5$- -- $0.5$- --
+Table 4: Subject conditioning improves generalisation on seen subjects but not novel subjects.
+We quote test accuracy at the point of best validation accuracy (early stopping).
+Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
+Table 4 reveals that in-distribution subject generalisation is helped by using subject conditioning methods.
+Both subject embeddings and FiLM conditioning outperform no conditioning, with FiLM being best.
+However, for novel subject generalisation, using a subject conditioning method makes little difference.
+Regardless of conditioning, we can generalise beyond chance-level, but suffer a reduction in accuracy compared to in-distribution subjects.
+Conditioning may be negligible in the unseen case as the embeddings are random rather than learned for the held-out subjects.
+Appendix C Aggregating Unlabelled Datasets
+To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets.
+As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS [47] and Cam-CAN [48, 53].
+We investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
+Balanced accuracy
+Labels Pre-training data (hours) Speech detection Voicing
+Armeni MOUS (160 hours) $0.5736\pm 0.0016$ $0.5220\pm 0.0013$
+Cam-CAN (159 hours) $\mathbf{0.5911}\pm 0.0009$ $\mathbf{0.5241}\pm 0.0004$
+Cam-CAN + MOUS (319 hours) $0.5872\pm 0.0012$ $0.5236\pm 0.0016$
+Gwilliams MOUS (160 hours) $0.5060\pm 0.0006$ $0.5175\pm 0.0008$
+Cam-CAN (159 hours) $\mathbf{0.5118}\pm 0.0011$ $\mathbf{0.5196}\pm 0.0010$
+Cam-CAN + MOUS (319 hours) $0.5116\pm 0.0017$ $0.5187\pm 0.0013$
+Table 5: Combining two unlabelled datasets does not improve over the best study.
+We quote test accuracy at the point of best validation accuracy (early stopping).
+The uncertainty is the standard error of the mean calculated over three seeds.
+The results in Table 5 suggest that combining these two datasets is not strictly better than the best dataset on its own.
+Pre-training on Cam-CAN alone leads to the best performance in all instances.
+This suggests that Cam-CAN is a better pre-training dataset than MOUS—a surprising result given that MOUS and Armeni et al.
+[2] share the same scanning hardware and MOUS includes data from subjects listening to speech while Cam-CAN has no such language tasks.
+This could be due to Cam-CAN being a cleaner dataset as we found artefacts can significantly affect learning.
+Combining Cam-CAN with MOUS performs marginally worse or the same as Cam-CAN alone.
+For all but one instance they are within the standard error of the mean of each other.
+The combined datasets always outperform MOUS on its own.
+In general, further investigation is necessary to determine how datasets should be best aggregated.
+Increasing the number of datasets could enable the network to model the variance between them and eventually improve over the best singular dataset.
+This remains an open question as we were only able to study up to two datasets in this paper.
+Appendix D Task Labels Are Hard To Scale; Generic SSL Scales Best
+Figure 5: Cross-task semi-supervised learning does not improve generalisation.
+We pre-train with increasing amounts of labelled data from Armeni et al.
+[2], using speech labels when evaluating voicing and voicing labels for speech.
+We also use a constant 38 hours of unlabelled data in pre-training.
+No pre-training shows test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
+without pre-training).
+SSL is when pre-training with only the unlabelled data.
+The shaded area shows the standard error of the mean across three seeds.
+Our self-supervised tasks led to clearly generalisable and scalable representations by providing an inductive bias towards speech decoding.
+However, could these representations be made more powerful by utilising labels from other speech decoding tasks to improve this bias?
+This moves us into the semi-supervised learning regime, where we use a mix of labelled and unlabelled data.
+With the same experimental setup as Figure 3, we add speech detection or voicing as an additional loss during pre-training.
+Figure 5 shows that no amount of additional labelled data improves downstream accuracy, with all data points remaining near the range of the standard error of the mean of the SSL baseline for both tasks.
+There also does not appear to be any indication of scaling within the range of labelled data available to us.
+This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy.
+The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space.
+Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks.
+This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech; nor does it mean more labelled data will not scale.
+However, the point remains salient.
+Labelled data is more difficult to scale given its scarcity and the generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
+Appendix E Experiment Details
 To generate the data in Table 1, we pre-train our representations with non-overlapping sample windows from all subjects and sessions 001 to 009 of Armeni et al.
@@
 For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting.
-We pre-train only with Shafto et al.
-[48], Taylor et al.
-[53].
+We pre-train only with Cam-CAN [48, 53].
 We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include.
@@
 Therefore, each speech detection label represents a 0.1 second period of time.
-Appendix C Hyperparameters
+Appendix F Hyperparameters
 We conduct a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture.
@@
 Table 9: Experimental hyperparameters.
-Appendix D Compute Resources
-All experiments were run on individual NVIDIA V100 GPUs with up to 32GiB of GPU memory on a system with up to 128GiB of RAM.
-Each pre-training run with the maximum amount of pre-training data took approximately 150 hours (6.25 days).
+Appendix G Compute Resources
+All experiments were run on individual NVIDIA V100 and A100 GPUs with up to 40GiB of GPU memory on a system with up to 1TiB of RAM.
+Each pre-training run with the maximum amount of pre-training data took approximately 200 hours (8.3 days).
 Fine-tuning following pre-training took up to another 12 hours.
-We estimate that we used approximately 2000 hours of compute for the final experimental runs, including hyperparameter searches.
+We estimate that we used approximately 3000 hours of compute for the final experimental runs, including hyperparameter searches.
 In total, over the course of developing this work from idea to final paper, we used around 10,000 hours of GPU compute.
-Appendix E Licences For Datasets And Code
+Appendix H Licences For Datasets And Code
 The Armeni et al.
@@
 [14] is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.
-Appendix F Broader Impacts
+Appendix I Broader Impacts
 Decoding speech from non-invasive brain recordings is likely to bring about significant positive societal impacts.

### v3

@@
 The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects.
-However, participants exhibit anatomical and other individual differences, and datasets use varied scanners and task designs.
+However, participants exhibit individual differences, such as anatomy, and datasets use varied scanners and task designs.
 As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets.
 In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning.
-To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings.
-Experimental results show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data.
+This gap exists for all neural data, but especially for magnetoencephalography (MEG), where the scale of individual datasets has not yet caught up with other modalities.
+To address this, we develop a set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings.
+Experimental results with MEG show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, outperform using the raw input representation, and even surpass comparable self-supervised approaches.
 In addition, we set new benchmarks for two foundational speech decoding tasks.
-Taken together, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
-Figure 1: Leveraging unlabelled data for speech decoding.
-We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled neuroimaging data, permitting learning from large heterogeneous datasets.
-The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation.
-We fine-tune the pre-trained network on labelled data, achieving faster generalisation owing to the strength of the representation learned in pre-training.
+Collectively, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
+Figure 1: Leveraging unlabelled data using pretext tasks for speech decoding.
+We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled MEG neuroimaging data, permitting learning from large heterogeneous datasets.
+The tasks apply a randomly selected neuroscientifically relevant transformation $T$ to the data and the network predicts the transformation.
+We then train a linear probe on top of the pre-trained model, which remains frozen, with labelled data, achieving superior generalisation (cf.
+raw inputs) owing to the strength of the representation.
 1 Introduction
-In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases [50].
-In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains [29, 7, 42, 46].
+In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases (Sutton, 2019).
+In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains (Jumper et al., 2021; Caron et al., 2021; OpenAI, 2023; Radford et al., 2023).
 In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated.
-State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech.
+State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech.
 Yet, a number of obstacles to scale remain.
-With few exceptions at present (e.g.
-[17]), speech decoding models tend not to train on data from more than one subject.
+With few exceptions at present, e.g.
+Défossez et al.
+(2023), speech decoding models tend not to train on data from more than one subject.
 Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks.
@@
 There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example.
-But overcoming these limitations, as has begun to happen in neighbouring sub-fields (e.g.
-[27]), holds the promise of training models on collective, internet-scale data.
-Given the scarcity of labelled data, self-supervised learning (SSL) appears promising as an avenue for domains where such data is rare or hard to obtain [3].
-In the SSL paradigm, pretext tasks pre-train a model on unlabelled data by generating implicit training labels through transformations of the input data in order to help a downstream task.
-We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals.
+But overcoming these limitations, as has begun to happen in neighbouring sub-fields, such as  Jiang et al.
+(2024), holds the promise of training models on collective, internet-scale data.
+While neuroimaging modalities such as electroencephalography (EEG) are more abundant, MEG may be a better modality for decoding as it provides a richer signal (Lopes da Silva, 2013; Hall et al., 2014).
+Given the scarcity of speech-labelled MEG data and the relative abundance of other MEG data, self-supervised learning (SSL) appears promising as it is an avenue for domains where labels are rare or hard to obtain (Balestriero et al., 2023).
+But the scale of public MEG data, while large, is still not at the volume of breakthroughs in self-supervised image and natural language processing, let alone EEG.
+Thus, SSL methods for MEG need to be highly data-efficient.
+Pretext tasks are one such method in which domain-specific self-supervised tasks are used to pre-train a model on unlabelled data by generating implicit training labels through transformations of the input in order to help a downstream task.
+We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals which we train using our pretext tasks.
 In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners.
-We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive magnetoencephalography (MEG), setting the baselines for speech detection and voicing classification on this data.
+We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive MEG, setting the baselines for speech detection and voicing classification on this data.
 The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time.
 Our main contributions are:
-  • A set of domain specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
-  • A neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
-  • A comprehensive experimental evaluation providing evidence that our approach can scale up with data and enable cross-dataset, task, and subject generalisation.
+  • A set of domain-specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
+  • A data-efficient neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
+  • A comprehensive experimental evaluation, using multiple times the volume of data in prior work, that verifies the above claims and additionally provides evidence for the existence of scaling laws when pre-training models with unlabelled MEG recordings.
 2 Related Work
 Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments.
-This is true both in recent state-of-the-art invasive studies [40, 39, 60, 9] and non-invasive studies [52].
+This is true both in recent state-of-the-art invasive studies (Moses et al., 2021; Metzger et al., 2023; Willett et al., 2023; Chen et al., 2024) and non-invasive studies (Tang et al., 2023).
 These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments.
-Focusing on semantic rather than phonetic decoding, work by Tang et al.
-[52] is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos.
+Focusing on semantic rather than phonetic decoding, the method developed by Tang et al.
+(2023) is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos.
 They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
@@
 Csaky et al.
-[11] find that a subset of individuals perform better when evaluated with a group-level model than with individual models.
+(2022) find that a subset of individuals perform better when evaluated with a group-level model than with individual models.
 Exploiting audio data in a multi-modal framework, Défossez et al.
-[17] show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated.
-Although they repeat the result within two MEG and two EEG datasets, Défossez et al.
-[17] do not show any improvements for pooling data across datasets.
-Moreover, they do not combine data from studies with different labels either; cf.
-[58, 16, 56].
-Unfortunately, two of these papers [58, 16] included a bug in their evaluation code.
-As such, their methods may perform no better than a baseline that provides pure noise inputs to the model [28].
+(2023) show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated.
+However, they do not demonstrate the ability to generalise to novel subjects and must retrain their model for new datasets.
+Moreover, although they repeat the result within two MEG and two EEG datasets, Défossez et al.
+(2023) do not show any improvements for pooling data across datasets.
+Their method is also unable to incorporate data without corresponding audio labels and so they do not combine data from studies with other kinds of labels either; cf.
+Wang & Ji (2022); Duan et al.
+(2023); Wang et al.
+(2023a).
+Unfortunately, the first two of these papers included a bug in their evaluation code.
+As such, their methods may perform no better than a baseline that provides pure noise inputs to the model (Jo et al., 2024).
 […]
-In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning [1, 15, 41, 31, 62, 21].
+In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning (Agrawal et al., 2015; Doersch et al., 2015; Noroozi & Favaro, 2016; Larsson et al., 2016; Zhang et al., 2016; Gidaris et al., 2018).
 Until now, similar approaches have not translated to the brain decoding literature.
@@
 For example, Jiang et al.
-[27] succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective.
-Although this work combined unlabelled EEG datasets, it only achieved improvements on non-speech tasks.
+(2024) succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective.
+Although this work combined unlabelled datasets, their results studied simpler non-speech tasks with EEG.
 Wang et al.
-[57] used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations.
-They attained impressive speech detection results but with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data.
+(2023b) used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations.
+Their impressive speech detection results were achieved with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data.
+Perhaps the closest work to ours in terms of unlocking scaling with neural data is BIOT (Yang et al., 2023).
+This is a self-supervised architecture for encoding bio-signals that is similarly capable of training with different datasets, labels, and varied numbers of sensors.
+Like the previous works, the approach tokenises signals for a transformer architecture, but instead of a masked loss it uses a contrastive pre-training objective.
+While theoretically supporting MEG, Yang et al.
+(2023) evaluate BIOT on simple ECG/EEG tasks rather than address the comparatively complex challenge of speech decoding with MEG data.
 3 Method
-To process continuous neuroimaging data, we introduce a neural architecture for encoding heterogeneous brain signals into latent representations.
-By developing pretext tasks with the objective of learning generalisable brain representations, we leverage this architecture for self-supervised learning from unlabelled data, hoping to replicate similar successes in computer vision [21, 8].
+To encode continuous neuroimaging data, we introduce a neural architecture to embed heterogeneous brain signals.
+We leverage this architecture for self-supervised learning from unlabelled MEG data using a set of pretext tasks designed to generate generalisable brain representations for speech decoding.
+With this approach, we hope to replicate similar successes in computer vision (Gidaris et al., 2018; Chen et al., 2020).
 3.1 Network Architecture
-We design a two-stage neural network architecture (Figure 2).
-The pre-training stage uses pretext tasks to train a representation with unlabelled brain data.
-Then, the fine-tuning stage uses this representation to learn the downstream task by fine-tuning with labelled data.
+Our two-stage neural network architecture (Figure 2) uses pretext tasks in pre-training to learn a representation with unlabelled brain data.
+Then, the fine-tuning stage uses this representation to learn the downstream task by training with labelled data.
 Figure 2: Architecture overview.
-Inputs are projected into a shared dimension, then encoded.
-In pre-training, all weights are trainable except for modules in light-red, while in shallow fine-tuning, modules with light-blue borders are frozen and modules with light-red borders are unfrozen.
-Deep fine-tuning is identical, except the encoder is trainable.
+Inputs are projected into a shared dimension by the dataset-conditional layer, then encoded.
+In pre-training, all weights are trainable except for modules in light-red, while in fine-tuning, modules with light-blue borders are frozen and modules with light-red borders are unfrozen.
 Dashed borders indicate optional components.
-Normalisation. We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance.
-Backbone.
+We divide recordings into windows of length $w$ seconds or $t$ samples.
+At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance.
 The network takes as input the standardised sample windows.
 To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$.
-Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs [61, 14].
-Specifically, our convolutional encoder adapts the implementation of the SEANet architecture [51] used in Défossez et al.
-[14].
+Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs (Zeghidour et al., 2022; Défossez et al., 2022).
+Specifically, our convolutional encoder adapts the implementation of the SEANet architecture (Tagliasacchi et al., 2020) used in Défossez et al.
+(2022) which we describe here and as part of Figure 2.
 As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions.
@@
 an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
-Pre-training.
+Just as speakers have different voices, neural responses between subjects have different characteristics.
+Consequently, individual variation leads to models that do not generalise well across subjects (Csaky et al., 2022).
+In the speech literature, models include speaker conditioning to account for these differences (Gibiansky et al., 2017).
+We take a similar approach by introducing subject conditioning.
+Zeghidour et al.
+(2022) find that conditioning is equally effective at the encoder bottleneck as in other stages of the model.
+Hence, we place ours at the cortex encoder bottleneck for simplicity.
+We use feature-wise linear modulation (FiLM) (Perez et al., 2018) as our conditioning method.
 Following the advice of Balestriero et al.
-[3, Section 3.2], we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation.
+(2023, Section 3.2), we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation.
 After the projector, linear classifiers make predictions for each of the pretext tasks.
-Fine-tuning.
-In this stage, we introduce classifiers for the downstream tasks and train with labelled data and supervised learning (Figure 2).
-Depending on the experiment, we fine-tune in one of two ways.
-In the first, we train an MLP classifier from scratch on top of the pre-trained representation, which remains frozen.
+When fine-tuning, we train a linear decoder, for a downstream task, on top of the pre-trained representation, which remains frozen.
 Thus, we backpropagate only through the classifier.
-We call this shallow fine-tuning.
-In the other, we fine-tune by training an MLP on top of the pre-trained representation again, but do not freeze the backbone, training end-to-end.
-Here, we backpropagate through the classifier and the backbone.
-We refer to this as deep fine-tuning.
-In both cases, a new trainable dataset-specific linear layer can be introduced for a novel dataset.
+A trainable dataset-specific linear layer can be introduced for a novel dataset.
 For speech detection, our classifier makes a prediction for each individual embedding.
 For voicing classification, where there is only one label for each sample window, the embeddings are flattened into a tensor in $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ representing the entire window.
-This is the input to the voicing classifier and is referred to as full epoch decoding in neuroimaging literature [12].
-Subject conditioning.
-Just as speakers have different voices, neural responses between subjects have different characteristics.
-Consequently, individual variation leads to models that do not generalise well across subjects [11].
-In the speech literature, models include speaker conditioning to account for these differences [20].
-We take a similar approach by introducing subject conditioning, represented as a subject embedding, to our model.
-With a SEANet-based architecture, Zeghidour et al.
-[61] find that conditioning is equally as effective at the encoder bottleneck as in other stages of the model.
-Hence, we place ours at the cortex encoder bottleneck for simplicity.
-We evaluate multiple types of conditioning including subject embeddings and feature-wise linear modulation (FiLM) [44].
+This is the input to the voicing classifier and is referred to as full epoch decoding in neuroimaging literature (Csaky et al., 2023).
 3.2 Pretext Tasks
-Our pretext tasks are unsupervised feature learning tasks for continuous neuroimaging signals that aim to learn generalisable speech decoding features.
-Since different datasets use different hardware and varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal.
-This means that these tasks do not require identifying specific sensor channels.
+Our pretext tasks are unsupervised feature learning tasks that aim to learn generalisable speech decoding features.
+Since different datasets use varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal.
 Band prediction.
-In the literature, neural responses can be broadly segmented into functional frequency bands [22, 45, 36].
-Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech [35], Theta ($\theta$) waves (4–8 Hz) reliably track [34] and phase-lock to the amplitude envelope of heard sentences [43], Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals [49], Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding [5] which affects lexical processing [59], Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g.
-memory, learning, reasoning, and planning) [19, 6], and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection [26] and phonemic feature classification in the STG [38] as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) [10].
+In the literature, neural responses can be segmented into functional frequency bands (Giraud & Poeppel, 2012; Piai et al., 2014; Mai et al., 2016).
+Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech (Luo et al., 2010), Theta ($\theta$) waves (4–8 Hz) reliably track (Luo & Poeppel, 2007) and phase-lock to the amplitude envelope of heard sentences (Peelle et al., 2012), Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals (Strauß et al., 2015), Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding (Bressler & Richter, 2015) which affects lexical processing (Weiss & Mueller, 2012), Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g.
+memory, learning, reasoning, and planning) (Fries, 2009; Buzsáki & Wang, 2012), and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection (Hamilton et al., 2018) and phonemic feature classification in the STG (Mesgarani et al., 2014) as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) (Cheung et al., 2016).
 As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
-To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected.
+To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding linear predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected.
 This yields the loss
@@
 Phase shift prediction.
-Phase coupling between networks of neuron populations is necessary for coordinating brain activity [18, 54].
+Phase coupling between networks of neuron populations is necessary for coordinating brain activity (Fries, 2005; Vidaurre et al., 2018).
 Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature.
-Supporting this insight, recent work [27] also finds phase to be an essential component of the signal.
+Supporting this insight, recent work (Jiang et al., 2024) also finds phase to be an essential component of the signal.
 To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho$ of the sensors.
@@
 Within speech, Hamilton et al.
-[26] find that localised regions of the STG respond to sustained speech and speech onsets.
+(2018) find that localised regions of the STG respond to sustained speech and speech onsets.
 Differentiating between neural responses from this region and others may be essential for decoding speech perception.
@@
 This includes understanding how well they can generalise across datasets, subjects, and tasks.
-We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG [32] and faster sampling rates than fMRI [25].
+We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG (Lopes da Silva, 2013) and faster sampling rates than fMRI (Hall et al., 2014).
+We pre-train all models to completion and then fine-tune on labelled data for each task.
+In all tables and figures, we quote the receiver operating characteristic area under the curve (ROC AUC) where chance is always 0.5 regardless of class balance.
+We show the test ROC AUC at the best validation ROC AUC (early stopping) and quote uncertainty as the standard error of the mean over three seeds.
+Additionally, we state the $t$-score and $p$-value from single-sample one-sided $t$-tests against chance.
 4.1 Experimental setup
 Datasets.
-Unless specified otherwise, our experiments use Cam-CAN [48, 53] as an unlabelled representation learning dataset for pre-training.
+Unless specified otherwise, our experiments use Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) as an unlabelled representation learning dataset for pre-training.
 This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings.
@@
 Armeni et al.
-[2] contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al.
-[24] has 27 subjects, each recorded for 2 hours (54 hours total).
+(2022) contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al.
+(2023) has 27 subjects, each recorded for 2 hours (54 hours total).
 Overall, we utilise over 200 hours of data.
@@
 Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise.
-Next, we downsample the signal to 250Hz, avoiding aliasing at frequencies up to our low-pass filter threshold.
+Treating the low-pass filter threshold as the Nyquist frequency, we downsample the signal to twice that at 250Hz, avoiding aliasing within our band of interest.
 Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
@@
 We select these tasks as they are simpler than phoneme recognition, but are foundational because they must be solved to decode speech accurately into natural language.
-Training.
-We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation.
-Appendix E provides complete training details for all experiments.
 4.2 Learning Generalisable Representations Using Pretext Tasks
-We investigate whether our self-supervised objectives produce useful representations by first learning a representation with each pretext task independently and then analysing its downstream generalisation using shallow fine-tuning (i.e.
-with the representation frozen).
-In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, and critically, outperform a randomly initialised and shallow fine-tuned baseline.
-The results are statistically significant in all cases, providing initial evidence that all of our tasks are helpful in learning representations that generalise downstream to speech decoding.
+Our first experiment investigates whether our self-supervised objectives produce generalisable representations.
+In Table 1, we show the results of pre-training models with each pretext task independently as well as together.
+Here, all of our pretext tasks lead to results that are statistically significant, and outperform a baseline fine-tuned without pre-training.
+This provides initial evidence that our tasks are helpful in speech decoding.
 Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own.
-As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding.
-For our other downstream task, voicing classification, we were not able to use shallow fine-tuning to generalise to a statistically significant degree with any of the pretext tasks nor with a random backbone.
-It may be too difficult to learn with shallow fine-tuning and so we turn to using deep fine-tuning for this task in the next section.
-Pretext task Speech detection balanced accuracy
-Armeni $t$ $p$ Gwilliams $t$ $p$
-Amp($\rho=0.2$) $0.5846\pm 0.0032$ $26$ $7\mathrm{e}{-4}$ $0.5116\pm 0.0004$ $29$ $6\mathrm{e}{-4}$
-Phase($\rho=0.5$) $0.5831\pm 0.0029$ $29$ $6\mathrm{e}{-4}$ $0.5093\pm 0.0014$ $7$ $1\mathrm{e}{-2}$
-Band $0.5941\pm 0.0024$ $39$ $2\mathrm{e}{-3}$ $0.5177\pm 0.0028$ $6$ $1\mathrm{e}{-2}$
-All $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
-None (rand. backbone) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
-Chance-level $0.5$- -- $0.5$- --
-Table 1: Pre-training with pretext tasks produces generalisable representations.
-We provide the test accuracy at the point of best validation accuracy (early stopping).
-In the none baseline, the network backbone is randomly initialised before fine-tuning (i.e.
-there is no pre-training).
+As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding than any individual task.
+Armeni Gwilliams
+Experiment ROC AUC $t$ $p$ ROC AUC $t$ $p$
+Linear $0.559\pm 2\mathrm{e}{-4}$ $341$ $4\mathrm{e}{-6}$ $0.527\pm 7\mathrm{e}{-5}$ $379$ $3\mathrm{e}{-6}$
+BIOT + linear $0.500\pm 4\mathrm{e}{-4}$ 00 $6\mathrm{e}{-1}$ $0.499\pm 2\mathrm{e}{-4}$ $-3$ $1\mathrm{e}{+0}$
+Ours No pre-training $0.519\pm 0.002$ $8$ $7\mathrm{e}{-3}$ $0.498\pm 0.003$ 00 $7\mathrm{e}{-1}$
++ linear Amp($\rho=0.2$) $0.602\pm 0.001$ $114$ $4\mathrm{e}{-5}$ $0.532\pm 0.005$ $6$ $1\mathrm{e}{-2}$
+Phase($\rho=0.5$) $0.603\pm 0.003$ $35$ $4\mathrm{e}{-4}$ $0.535\pm 0.003$ $12$ $3\mathrm{e}{-3}$
+Band $0.616\pm 0.003$ $44$ $3\mathrm{e}{-4}$ $0.542\pm 0.001$ $46$ $2\mathrm{e}{-4}$
+All tasks $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $\mathbf{0.543}\pm 0.003$ $13$ $3\mathrm{e}{-3}$
+Table 1: Pre-training with pretext tasks leads to better representations for decoding speech.
+In the linear-only case, we train a supervised linear classifier on the input MEG signals.
+For BIOT, we train a linear layer on top of the pre-trained BIOT representation, with the rest of the model frozen.
+Similarly, for ours, we train a linear probe on top of our pre-trained backbone with its weights frozen.
+In the no pre-training baseline, the backbone uses randomly initialised and subsequently unmodified weights.
 When all pretext tasks are used, their losses are weighted equally.
-The uncertainty is the standard error of the mean over three random seeds.
-We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
-Models are pre-trained on all of Cam-CAN [48, 53] and fine-tuned on Armeni et al.
-[2] or Gwilliams et al.
-[24] respectively.
-Among the pretext tasks, band prediction performs best.
-This could be because the representation easily identifies phase-locking to speech onset in theta waves [43].
-However, further investigation is necessary to determine which frequency band is most informative.
-The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix F).
+Now, we turn to the other baselines.
+Our approach significantly outperforms the equivalent with a raw MEG input instead of a pre-trained representation (the linear experiment).
+Here, the baseline has substantially more trainable parameters because the input dimension is far larger without an encoder.
+Even with this bias favouring the experiment with the raw input, using our representation still performs better.
+We also compare our approach to BIOT (Yang et al., 2023) which is a similar state-of-the-art self-supervised method.
+When BIOT is pre-trained using exactly the same data, the fine-tuned probe fails to generalise entirely after exhaustive hyperparameter tuning.
+We put this down to three critical reasons.
+Firstly, BIOT was designed around considerably lower-dimensional signals.
+Their EEG evaluation used an order of magnitude fewer sensors than our MEG data.
+With MEG, their transformer approach requires many more channel embeddings, leading to difficulty learning the complex interactions between sensors.
+Secondly, our self-supervised objective extracts speech decoding features which is essential for solving speech decoding tasks.
+BIOT performs well on simple EEG tasks in Yang et al.
+(2023)’s evaluation, but non-invasive speech decoding is significantly more challenging.
+Together, these obstacles suggest a vast amount of data is required to learn their objective with MEG.
+Indeed, given that they pre-train with over 50 thousand hours of EEG data in their evaluation, their objective appears too general to efficiently learn a representation for speech decoding from the limited amount of MEG pre-training data (160 hours) available to us.
+This highlights the importance of data-efficiency in SSL methods for MEG.
+Among the individual pretext tasks, band prediction leads the rest.
+Perhaps this is because, by learning to discriminate between meaningful bands, the representation easily identifies phase-locking to speech onset in theta waves (Peelle et al., 2012).
+Further investigation is necessary here.
+The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search.
 We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors.
-Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset [26].
-In contrast, a larger $\rho$ for phase shift prediction may be best because phase is descriptive of neural synchrony which is distributed across the brain.
-Unlike localised responses, a large proportion of the sensors can detect this feature.
-Whilst we have shown that our representations generalise, do they outperform supervised baselines when trained to saturation?
-Using shallow fine-tuning once more, we compare our approach directly to similarly parameterised supervised classifiers in Table 2.
-A representation pre-trained with all pretext tasks significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets.
-Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is far larger without an encoder.
-Even with this bias favouring the supervised classifier, the self-supervised case performs better.
-Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
-Model Speech detection balanced accuracy
-Armeni $t$ $p$ Gwilliams $t$ $p$
-Supervised
-Linear $0.5222\pm 0.0015$ $12$ $3\mathrm{e}{-3}$ $0.5006\pm 0.0001$ $5$ $2\mathrm{e}{-2}$
-Self-supervised + fine-tuning
-Random backbone
-   MLP (two layers) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
-Pre-trained backbone
-   Linear $0.5853\pm 0.0017$ $52$ $3\mathrm{e}{-4}$ $0.5011\pm 0.0001$ $9$ $6\mathrm{e}{-3}$
-   MLP (two layers) $\mathbf{0.6011}\pm 0.0018$ $56$ $2\mathrm{e}{-4}$ $\mathbf{0.5206}\pm 0.0010$ $22$ $1\mathrm{e}{-3}$
-Chance-level $0.5$- -- $0.5$- --
-Table 2: Self-supervised pre-training outperforms supervised classifiers.
-We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion.
-We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds.
-We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance.
-Self-supervised models are pre-trained on all of Cam-CAN [48, 53] and shallow fine-tuned on Armeni et al.
-[2] or Gwilliams et al.
-[24] while the supervised models are trained entirely on the latter datasets.
+Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018).
+In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised.
+As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.
 4.3 Scaling Speech Decoding With Unlabelled Data
-Figure 3: Scaling unlabelled data improves downstream generalisation.
-We pre-train the model on increasing amounts of unlabelled data from Cam-CAN [48, 53] and fine-tune on voicing and speech detection on the Armeni et al.
-[2] and Gwilliams et al.
-[24] datasets.
-No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
-without pre-training).
-We show only one no pre-training line as the result is the same for both datasets.
-The shaded area shows the standard error of the mean across three seeds.
-Previously, we showed that the combination of all pretext tasks leads to the best representations and that this outperforms a randomly initialised backbone as well as a supervised classifier.
-In this section, we examine how these pretext tasks fare as we increase the amount of unlabelled data, analysing scaling performance on downstream tasks.
-As before, we pre-train with the combined pretext tasks using Cam-CAN [48, 53], which does not have any speech decoding labels, and now fine-tune for ten epochs on Armeni et al.
-[2] and Gwilliams et al.
-[24], showcasing transfer with minimal fine-tuning.
-This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI.
-Where earlier we analysed only speech detection, here, we also focus on voicing classification by using deep fine-tuning for this task only.
-As evidenced by Section 4.2, we require this because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
-Figure 3 shows balanced accuracy as we increase the amount of unlabelled data in pre-training up to approximately 160 hours.
-For both tasks, pre-training outperforms the no pre-training baseline, which fails to generalise at all, showing that any unlabelled data is sufficient to pre-train a useful representation.
-In both tasks, there is a clear improvement in accuracy as the amount of unlabelled data increases.
-Thus, adding unlabelled data improves generalisation and scales performance.
-With speech detection, the improvement in accuracy is more significant for Armeni et al.
-[2].
-Noting that there is more within-subject data in this dataset than Gwilliams et al.
-[24], speech detection may be more sensitive to subject variation than voicing classification.
-We scaled up the pre-training dataset by increasing the number of subjects.
-Since this led to consistent and almost monotonic improvements in downstream accuracy, our self-supervised method is an exception to the common consensus that pooling subjects worsens generalisation.
-Moreover, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation.
-This is particularly notable as the Armeni et al.
-[2], Gwilliams et al.
-[24], and pre-training datasets all use different scanners entirely.
+Figure 3: Scaling unlabelled data improves generalisation.
+We pre-train the model on increasing amounts of unlabelled data from Cam-CAN (Shafto et al., 2014; Taylor et al., 2017).
+The solid lines are the best linear fits to the data and the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
+Here, we analyse generalisation as we increase the volume of unlabelled data, analysing scaling performance on downstream tasks.
+As before, we pre-train with the combined pretext tasks.
+Figure 3 shows ROC AUC as we increase the amount of unlabelled data in pre-training up to approximately 160 hours.
+For both tasks, pre-training with any amount of data is sufficient to beat chance and there is a clear improvement in accuracy as the amount of unlabelled data increases.
+For speech detection on Armeni et al.
+(2022), scaling appears logarithmic in log-space; for all others, ROC AUC improves log-linearly within the data regime we study.
+In any case, adding unlabelled data has improved generalisation.
+Notably, we have scaled far beyond the data regime of prior surgical and non-surgical work and yet performance has continued to scale.
+Thus, our self-supervision approach may remain useful as the volume of open data in the field continues to rapidly increase.
+Our results also reveal several new and notable phenomena.
+Firstly, we scaled up the pre-training dataset by increasing the number of subjects.
+Since this led to consistent and almost monotonic improvements in downstream accuracy, our method is an exception to the common consensus that pooling subjects worsens generalisation.
+Secondly, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation.
+This is particularly surprising as the Armeni et al.
+(2022), Gwilliams et al.
+(2023), and our pre-training dataset all use different scanners entirely.
 Performing well across these datasets indicates that, together, our architecture and pretext tasks successfully generate representations that are generalisable across heterogeneous scanners.
-Beyond hardware, our results also show improvements on distinct tasks, namely speech detection and voicing classification, thus illustrating cross-task generalisation.
-Thus, our pretext tasks are sufficiently generic to produce representations that generalise to multiple downstream tasks, indicating their applicability to speech decoding more generally.
-Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks.
-Remarkably, this shows that brain data collected from any task (including those that are not linguistic) can be used to scale downstream speech decoding performance.
-4.4 Scaling Unlabelled Data Improves Zero-Shot Generalisation To Novel Subjects
-Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11].
-Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge.
+Finally, we note that our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks.
+Remarkably, this shows that unlabelled brain data collected from any task (including those that are not linguistic) can be used to improve speech decoding performance.
+Since the results show improvements on both downstream tasks, this indicates that our pretext tasks are sufficiently generic to produce representations that work with multiple speech decoding tasks while still generalising well on each task individually.
+This is generally a challenging trade-off to manage.
+However, we notice that in both tasks, the base accuracy is higher and the improvement in ROC AUC is steeper for Armeni et al.
+(2022).
+This is likely to be because this dataset has more within-subject data.
+The weaker results for Gwilliams et al.
+(2023) may be a consequence of the larger number of subjects with shorter intra-subject recordings and greater subject variation.
+These observations support the findings of other recent work such as Csaky et al.
+(2022).
+4.4 Scaling Unlabelled Data Improves Generalisation To Novel Subjects
+In neuroimaging, brain data is generally highly variable across participants, leading to difficulty transferring models to novel subjects (Csaky et al., 2022).
+Whilst we have shown generalisation across subjects, here, we investigate whether we can generalise to novel subjects—an even more difficult challenge.
 This is critical in order to widely deploy speech BCIs for new patients.
-In the following experiment, we pre-train using the same setup as Figure 3, but fine-tune only on Gwilliams et al.
-[24].
-We use FiLM conditioning as this performs best for seen and unseen subjects (Appendix B).
-During training, we hold out three subjects with which we evaluate novel subject generalisation (i.e.
-unseen subject performance).
-Figure 4: Scaling pre-training data improves novel subject generalisation.
-As before, we pre-train with increasing amounts of unlabelled data from Cam-CAN [48, 53], but now fine-tune only on Gwilliams et al.
-[24].
-In the seen case, we evaluate on held-out sessions from subjects that were used in the training data.
-For the unseen case we evaluate on three completely held-out subjects.
-No pre-training shows the best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
-without pre-training).
-We show only one no pre-training line as the result is the same for both datasets.
-The shaded area is the standard error of the mean over three seeds.
-Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive trend in performance for unseen (novel) subjects.
-This is a promising result, indicating that scaling with our method is an encouraging direction for resolving the issues related to individual variance faced by prior work.
+In this experiment, we fine-tune only on Gwilliams et al.
+(2023) and hold out three subjects with which we evaluate novel subject generalisation.
+Figure 4: Scaling unlabelled data improves novel subject generalisation.
+We fine-tune on Gwilliams et al.
+(2023).
+When in-distribution, we evaluate on held-out sessions from subjects in the training set; when out-of-distribution, we evaluate on three held-out subjects.
+The solid lines are the best linear fits while the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
+Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive log-linear trend in performance for novel subjects.
+This indicates that scaling our method is an encouraging direction for resolving the challenges of subject variance faced by prior work.
+Moreover, as far as we are aware, this is the first result to demonstrate novel subject generalisation in speech decoding from MEG.
+4.5 Aggregating Unlabelled MEG Datasets
+To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets.
+As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS (Schoffelen et al., 2019) and Cam-CAN (Shafto et al., 2014; Taylor et al., 2017).
+In this section, we investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
+Pre-training Armeni Gwilliams
+dataset Hours ROC AUC $t$ $p$ ROC AUC $t$ $p$
+Cam-CAN 159 $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $0.543\pm 0.003$ $13$ $3\mathrm{e}{-3}$
+MOUS 160 $0.605\pm 0.000$ $261$ $7\mathrm{e}{-6}$ $0.543\pm 0.004$ $9$ $5\mathrm{e}{-3}$
+Cam-CAN + MOUS 319 $0.611\pm 0.003$ $40$ $3\mathrm{e}{-4}$ $\mathbf{0.546}\pm 0.002$ $20$ $1\mathrm{e}{-3}$
+Table 2: Combining unlabelled datasets shows signs of outperforming single studies.
+We examine performance on the speech detection task.
+We see a small improvement when the datasets are combined on Gwilliams et al.
+(2023), but not Armeni et al.
+(2022).
+The results in Table 2 show, for the first time, that combining datasets can improve performance on downstream speech decoding tasks.
+It leads to better performance on Gwilliams et al.
+(2023) compared to pre-training on either dataset alone.
+Interestingly, this was not the case for Armeni et al.
+(2022) where pre-training on Cam-CAN alone performed best.
+Combined pre-training did, however, outperform training only on MOUS.
+It is surprising that pre-training on Cam-CAN was better than pre-training on MOUS when evaluating on Armeni et al.
+(2022) given that MOUS and Armeni et al.
+(2022) both used speech tasks and were acquired on the same MEG scanner.
+Cam-CAN, by contrast, did not use a speech task and was acquired on a different MEG scanner.
+We hypothesise that the better results for Cam-CAN are due to it being a cleaner dataset.
+During our experiments, we found that data quality, even among unlabelled data, can have a significant affect as artefacts in recordings disrupt learning.
+While the combination of the two datasets includes far more hours of data than any prior work on deep learning with MEG, further work needs to be done to aggregate more datasets.
+Here, we were limited by compute budget.
+Increasing the number of datasets could enable the network to eventually always improve over the best singular dataset.
+Just as increasing the number of subjects (rather than only within-subject data) improves novel subject generalisation, a larger number of datasets may be key to scaling results when datasets are aggregated in pre-training.
 4.6 Limitations
@@
 brain-to-text).
-This has been achieved with surgical data [40, 60] but not yet convincingly with non-invasive methods like MEG or EEG [28].
-Speech detection has played an important role in the development of full brain-to-text in a surgical context [40] and we hope may play a similar role for non-invasive methods.
-Prior work has further used voice classification as a stand in for phoneme classification [23], and we have been able to improve on these results here.
+This has been achieved with surgical data (Moses et al., 2021; Willett et al., 2023) but not yet convincingly with non-invasive methods like MEG or EEG (Jo et al., 2024).
+Speech detection has played an important role in the development of full brain-to-text in a surgical context (Moses et al., 2021) and we hope may play a similar role for non-invasive methods.
+Prior work has further used voice classification as a stand in for phoneme classification (Gwilliams et al., 2022), and we have been able to improve on these results here.
 In future work, we would like to expand this to all English phonemes.
-Secondly, while we have been able to demonstrate the utility of a few pretext tasks here, we do not claim to have exhausted the full set of useful tasks.
+Secondly, while we have been able to demonstrate the utility of a few pretext tasks, we do not claim to have exhausted the full set of useful tasks.
 Rather, we conjecture that more useful pretext tasks remain to be found and believe a useful avenue of research will be into other input representations for brain recordings.
@@
 Ultimately, solving speech decoding could transform the lives of patients with severe communication difficulties.
-Yet, this promise has not materialised because the field has been blocked by its inability to scale up data to leverage deep learning.
+This promise has not yet materialised because the field has been blocked by its inability to scale up data to leverage deep learning.
 Prior methods have been unable to aggregate data across different datasets, labels, or subjects to scale up because of heterogeneity in recording hardware, experiment design, and participants.
@@
 But until now, no one has developed a general solution.
-We provided a unified method that leverages unlabelled data using generic pretext tasks that shows that all of these problems can be solved.
+We provided a unified method that leverages unlabelled recordings data-efficiently using generic pretext tasks that shows that all of these problems can be solved.
 We verified this with experiments showing that our method not only scales with heterogeneous data but even generalises across datasets, subjects, and tasks.
@@
 If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
-Acknowledgments and Disclosure of Funding
-We would like to thank Botos Csaba for many insightful discussions and creative ideas which helped shaped the direction of this work.
-Thanks also to Minqi Jiang for an encouraging conversation on unsupervised representation learning, Mats W.J.
-van Es for technical assistance with the OSL library, Brian Liu for technical contributions which did not reach the final paper, and Miran Özdogan for reviewing a draft of this work.
-The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work.
-http://dx.doi.org/10.5281/zenodo.22558.
-[…]
-Appendix A Supervised Heard Speech Decoding
-While this work is primarily about scaling speech decoding, we also set the first baselines for speech detection and voicing classification from MEG with connected auditory speech stimuli.
-Therefore, in this section we provide fully supervised results where our architecture is trained to completion using all available labelled data in the corresponding dataset (except the validation and test splits) and compare this to a linear classifier baseline.
-These results are not comparable with our self-supervised models because the entire architecture is trained end-to-end.
-In our SSL work with shallow fine-tuning, only the final classifier is trained with labels.
-Experiment Balanced accuracy
-Speech detection Voicing
-Armeni (linear) $0.5222\pm 0.0015$ $0.5207\pm 0.0007$
-Armeni (ours) $\mathbf{0.7253}\pm 0.0005$ $\mathbf{0.5282}\pm 0.0006$
-Gwilliams (linear) $0.5006\pm 0.0001$ $0.5166\pm 0.0002$
-Gwilliams (ours) $\mathbf{0.5946}\pm 0.0016$ $\mathbf{0.5211}\pm 0.0005$
-Table 3: Supervised baselines. Uncertainty is standard error of the mean over three random seeds.
-The results in Table 3 show that we can achieve strong speech detection results on Armeni et al.
-[2] and Gwilliams et al.
-[24] with our architecture.
-The difference between the two may be due to a combination of different scanners, numbers of sensors (208 vs 269 magnetometers), varied difficulty in the audio stimulus, and more within-subject data in Armeni et al.
-[2].
-We also achieve good results on voicing classification for both datasets, easily outperforming the linear models.
-One potential way to improve on both supervised and self-supervised approaches is by using a temporally aware model such as an LSTM or an attentional model such as a Transformer as the classifier.
-We do not investigate this within this paper as it is out of scope.
-Rather, we choose to use simple classifiers as the purpose of this paper is to demonstrate the potential of scaling.
-Appendix B Subject Conditioning
-To investigate subject conditioning approaches, we pre-train then fine-tune our models on a subset of Gwilliams et al.
-[24], holding out three subjects for evaluation.
-We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation.
-We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation.
-For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
-Subject Subject voicing balanced accuracy
-conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
-None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
-Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
-FiLM [44] $\mathbf{0.5197}\pm 0.001$ $18.6$ $0.001$ $\mathbf{0.5058}\pm 0.002$ $3.3$ $0.04$
-Chance-level $0.5$- -- $0.5$- --
-Table 4: Subject conditioning improves generalisation on seen subjects but not novel subjects.
-We quote test accuracy at the point of best validation accuracy (early stopping).
-Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
-Table 4 reveals that in-distribution subject generalisation is helped by using subject conditioning methods.
-Both subject embeddings and FiLM conditioning outperform no conditioning, with FiLM being best.
-However, for novel subject generalisation, using a subject conditioning method makes little difference.
-Regardless of conditioning, we can generalise beyond chance-level, but suffer a reduction in accuracy compared to in-distribution subjects.
-Conditioning may be negligible in the unseen case as the embeddings are random rather than learned for the held-out subjects.
-Appendix C Aggregating Unlabelled Datasets
-To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets.
-As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS [47] and Cam-CAN [48, 53].
-We investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
-Balanced accuracy
-Labels Pre-training data (hours) Speech detection Voicing
-Armeni MOUS (160 hours) $0.5736\pm 0.0016$ $0.5220\pm 0.0013$
-Cam-CAN (159 hours) $\mathbf{0.5911}\pm 0.0009$ $\mathbf{0.5241}\pm 0.0004$
-Cam-CAN + MOUS (319 hours) $0.5872\pm 0.0012$ $0.5236\pm 0.0016$
-Gwilliams MOUS (160 hours) $0.5060\pm 0.0006$ $0.5175\pm 0.0008$
-Cam-CAN (159 hours) $\mathbf{0.5118}\pm 0.0011$ $\mathbf{0.5196}\pm 0.0010$
-Cam-CAN + MOUS (319 hours) $0.5116\pm 0.0017$ $0.5187\pm 0.0013$
-Table 5: Combining two unlabelled datasets does not improve over the best study.
-We quote test accuracy at the point of best validation accuracy (early stopping).
-The uncertainty is the standard error of the mean calculated over three seeds.
-The results in Table 5 suggest that combining these two datasets is not strictly better than the best dataset on its own.
-Pre-training on Cam-CAN alone leads to the best performance in all instances.
-This suggests that Cam-CAN is a better pre-training dataset than MOUS—a surprising result given that MOUS and Armeni et al.
-[2] share the same scanning hardware and MOUS includes data from subjects listening to speech while Cam-CAN has no such language tasks.
-This could be due to Cam-CAN being a cleaner dataset as we found artefacts can significantly affect learning.
-Combining Cam-CAN with MOUS performs marginally worse or the same as Cam-CAN alone.
-For all but one instance they are within the standard error of the mean of each other.
-The combined datasets always outperform MOUS on its own.
-In general, further investigation is necessary to determine how datasets should be best aggregated.
-Increasing the number of datasets could enable the network to model the variance between them and eventually improve over the best singular dataset.
-This remains an open question as we were only able to study up to two datasets in this paper.
-Appendix D Task Labels Are Hard To Scale; Generic SSL Scales Best
-Figure 5: Cross-task semi-supervised learning does not improve generalisation.
-We pre-train with increasing amounts of labelled data from Armeni et al.
-[2], using speech labels when evaluating voicing and voicing labels for speech.
-We also use a constant 38 hours of unlabelled data in pre-training.
-No pre-training shows test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e.
-without pre-training).
-SSL is when pre-training with only the unlabelled data.
-The shaded area shows the standard error of the mean across three seeds.
-Our self-supervised tasks led to clearly generalisable and scalable representations by providing an inductive bias towards speech decoding.
-However, could these representations be made more powerful by utilising labels from other speech decoding tasks to improve this bias?
-This moves us into the semi-supervised learning regime, where we use a mix of labelled and unlabelled data.
-With the same experimental setup as Figure 3, we add speech detection or voicing as an additional loss during pre-training.
-Figure 5 shows that no amount of additional labelled data improves downstream accuracy, with all data points remaining near the range of the standard error of the mean of the SSL baseline for both tasks.
-There also does not appear to be any indication of scaling within the range of labelled data available to us.
-This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy.
-The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space.
-Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks.
-This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech; nor does it mean more labelled data will not scale.
-However, the point remains salient.
-Labelled data is more difficult to scale given its scarcity and the generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
-Appendix E Experiment Details
-To generate the data in Table 1, we pre-train our representations with non-overlapping sample windows from all subjects and sessions 001 to 009 of Armeni et al.
-[2] for 200 epochs with a $0.8:0.1:0.1$ train/val/test split.
-We shallow fine-tune with labels for 30 epochs using all subjects, but only session 010, which was held-out during pre-training.
-For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting.
-We pre-train only with Cam-CAN [48, 53].
-We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include.
+Ethics Statement
+In this work, we use data from studies that involve human subjects (Armeni et al., 2022; Gwilliams et al., 2023; Shafto et al., 2014; Taylor et al., 2017; Schoffelen et al., 2019).
+These datasets are public, cited, and have their own ethical approvals.
+The documentation for these is available with the publications for the respective datasets.
+While there are clear positive impacts, we acknowledge that insights from neural speech decoding research may not all be beneficial.
+Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g.
+stammering).
+As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants.
+Nevertheless, the maturity of this technology could also present potential negative societal impacts.
+For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals.
+Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent.
+Moreover, there are possible long horizon effects associated with speech decoding research.
+Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society.
+In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality.
+Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse.
+Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.
+Reproducibility Statement
+In the supplementary materials, we have provided an anonymised code repository with instructions for reproducing our main experiments.
+We also include details on experiment design and setup (Section 4.1 and Appendix A), hyperparameters (Appendix B), and compute (Appendix C).
+While we attempt to be exhaustive with these details, any information not found directly in the main body or appendices can be located in the supplementary materials.
+Appendix A Experiment Details
+We pre-train with non-overlapping sample windows from all subjects and sessions.
+We adjust the amount of unlabelled data used from Cam-CAN by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include.
 Each seed uses a different set of subjects to reduce negative effects from outlier subjects.
-We fine-tune with all data in Armeni et al.
-[2] and Gwilliams et al.
-[24] with the same train/val/test ratios as before.
-For the results shown in Figure 5, we use the same setup as in Figure 3.
-In addition to the unlabelled data, we add labelled data from Armeni et al.
-[2], increasing the number of sessions used in the sequence 1, 3, 6, 9, 12, and 27.
-For this experiment, we calculate the number of labelled data hours using an estimate of 1 hour per session per subject.
-Finally, for the results in Table 4, we pre-train and deep fine-tune to saturation with subjects 01-07 of Gwilliams et al.
-[24] and test zero-shot generalisation using held-out subjects 08-10.
-All other experimental details are the same as the previous two experiments.
-In all experiments, we use three randomly selected seeds for each pre-training and corresponding fine-tuning run.
+When fine-tuning with Armeni et al.
+(2022), we hold out session 010 from all subjects during training and validation, using this for evaluation.
+Similarly, when fine-tuning with Gwilliams et al.
+(2023), we hold out session 1 from subjects 23, 24, 25, 26, and 27, using these sessions for evaluation only.
+As there is limited within-subject data in the latter dataset, we did not hold out a session from all subjects as before.
+For our novel subject experiments, we hold out subjects 1, 2, and 3 entirely and use the data for these subjects during evaluation.
+In Gwilliams et al.
+(2023), we note that they use four different tasks for each subject and their order is randomized between subjects.
+Both sessions for each task are repeats of the task.
+This means that while the recording itself is unseen, in this dataset, it is possible that heldout sessions use tasks that may have been seen in the training set.
+In all experiments, we always fine-tune to completion (usually around 300 epochs), taking the test metric at the best validation loss (early stopping).
+We use three randomly selected seeds for each pre-training and corresponding fine-tuning run.
 For speech detection, since our encoder reduces the temporal dimension from 125 samples (the number of samples in a 0.5 second window with a sample rate of 250Hz) down to 5 embeddings, we downsample our speech detection labels to match using PyTorch’s torch.nn.functional.interpolate.
 Therefore, each speech detection label represents a 0.1 second period of time.
-Appendix F Hyperparameters
-We conduct a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture.
-In all experiments in this section, we measure downstream performance (using shallow fine-tuning) with best validation accuracy for speech detection unless explicitly specified, training and evaluating only on data from subject 001 of Armeni et al.
-[2] to minimise training time, over three random seeds.
-All uncertainties in this section are standard deviations and not standard error of the mean as in the main body.
-Table 6 shows the results of a search over the proportion $\rho$ of sensors to apply the phase and amplitude pretext task transformations to.
-We only search up to $\rho=0.5$ as anything above this is equivalent to applying the inverse transformation to a proportion $1-\rho$ of the sensors.
-We find that for phase shift prediction, $\rho=0.5$ is optimal, while for amplitude scale prediction $\rho=0.2$ is best.
-Task $\rho$ Balanced accuracy
-Phase shift prediction 0.1 $0.5452\pm 0.0048$
-0.2 $0.5566\pm 0.0014$
-0.3 $0.5578\pm 0.0021$
-0.4 $0.5530\pm 0.0014$
-0.5 $\mathbf{0.5655}\pm 0.0026$
-Amplitude scale prediction 0.1 $0.5440\pm 0.0012$
-0.2 $\mathbf{0.5796}\pm 0.0050$
-0.3 $0.5625\pm 0.0063$
-0.4 $0.5684\pm 0.0016$
-0.5 $0.5650\pm 0.0026$
-Table 6: Best $p$ for phase and amplitude pretext tasks.
-In our architecture, we define $d_{\mathrm{backbone}}$ to be the bottleneck dimension of the cortex encoder (i.e.
-the embedding dimension) as well as the size of $d_{\mathrm{shared}}$.
-Balestriero et al.
-[3], using the results from work by Bordes et al.
-[4], suggest that increasing the backbone output dimension in self-supervised learning leads to better downstream transfer.
-We investigate this for our architecture by varying the encoder dimension in Table 7.
-Our results are consistent with their findings.
-Balanced accuracy is highest at the largest encoder dimension we evaluated.
-Encoder dimensions larger than $2048$ led to unstable training and were not feasible given the compute resources available to us.
-$d_{\mathrm{backbone}}$ Balanced accuracy
-$128$ $0.5775\pm 0.0037$
-$300$ $0.5715\pm 0.0012$
-$512$ $0.5835\pm 0.0016$
-$1024$ $0.5800\pm 0.0050$
-$2048$ $\mathbf{0.5987}\pm 0.0024$
-Table 7: Optimal encoder dimension.
-In Table 8, we show the effect of varying the length $w$, i.e.
-the window size of each data sample.
-Increasing the window to five seconds led to unstable losses in pre-training so we did not evaluate above a window length of two seconds.
-This is likely to be because increasing the window size reduces the number of training samples when only non-overlapping samples are used.
-For speech detection, there appears to be a U-shaped distribution, where very short and very long windows lead to better downstream accuracy.
-For voicing classification, the window lengths we evaluated seem to make little difference.
-However, windows smaller than 0.1 seconds may lead to significantly lower accuracy as most phonemes will occur for at least this length of time in the neural response [23].
-Window length (s) Balanced accuracy
-Speech detection Voicing classification
-$0.2$ $0.5919\pm 0.0014$ $0.5033\pm 0.0014$
-$0.4$ $0.5804\pm 0.0036$ $\mathbf{0.5077}\pm 0.0027$
-$0.5$ $0.5715\pm 0.0012$ $0.5055\pm 0.0023$
-$1.0$ $0.5790\pm 0.0012$ $0.5062\pm 0.0020$
-$2.0$ $\mathbf{0.6116}\pm 0.0019$ $0.5065\pm 0.0011$
-Table 8: Ideal sample window length.
-We also evaluated several different configurations of the SEANet architecture for the cortex encoder, converging on convolutional blocks with channel dimensions $(512,512,512,512)$.
-While these ablations indicate a theoretically ideal architectural configuration, in practice, we altered our final experimental architecture due to instabilities during training when data was scaled up.
+Appendix B Hyperparameters
+We conducted a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture.
+While these ablations indicated a theoretically ideal architectural configuration, in practice, we altered our final experimental architecture due to instabilities during training when data was scaled up.
 Our final architecture hyperparameters achieve a balance between the best values from our hyperparameter search and stable training.
@@
 Pre-training epochs $200$
-Optimizer AdamW [33]
+Optimizer AdamW (Loshchilov & Hutter, 2019)
 Learning rate $0.000066$
@@
 Table 3: Experimental hyperparameters.
-Appendix G Compute Resources
+Appendix C Compute Resources
 All experiments were run on individual NVIDIA V100 and A100 GPUs with up to 40GiB of GPU memory on a system with up to 1TiB of RAM.
@@
 In total, over the course of developing this work from idea to final paper, we used around 10,000 hours of GPU compute.
-Appendix H Licences For Datasets And Code
+Appendix D Licences For Datasets And Code
 The Armeni et al.
-[2] dataset is distributed under CC-BY-4.0 while the Gwilliams et al.
-[24] dataset is distributed under the CC0 1.0 Universal licence.
+(2022) dataset is distributed under CC-BY-4.0 while the Gwilliams et al.
+(2023) dataset is distributed under the CC0 1.0 Universal licence.
 The Schoffelen et al.
-[47] dataset is distributed with a RU-DI-HD-1.0 licence from the Donders institute.
-The licence for the Cam-CAN [48, 53] dataset is unknown.
+(2019) dataset is distributed with a RU-DI-HD-1.0 licence from the Donders institute.
+The licence for the Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) dataset is unknown.
 The SEANet code adapted from Défossez et al.
-[14] is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.
-Appendix I Broader Impacts
-Decoding speech from non-invasive brain recordings is likely to bring about significant positive societal impacts.
-Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g.
-stammering).
-As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants.
-Nevertheless, the maturity of this technology could also present potential negative societal impacts.
-For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals.
-Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent.
-Moreover, there are possible long horizon effects associated with speech decoding research.
-Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society.
-In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality.
-Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse.
-Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.
+(2022) is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.

### v4

@@
 Abstract
-The past few years have produced a series of spectacular advances in the decoding of speech from brain activity.
-The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects.
-However, participants exhibit individual differences, such as anatomy, and datasets use varied scanners and task designs.
-As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets.
-In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning.
-This gap exists for all neural data, but especially for magnetoencephalography (MEG), where the scale of individual datasets has not yet caught up with other modalities.
-To address this, we develop a set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings.
-Experimental results with MEG show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, outperform using the raw input representation, and even surpass comparable self-supervised approaches.
-In addition, we set new benchmarks for two foundational speech decoding tasks.
-Collectively, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
+The past few years have seen remarkable progress in the decoding of speech from brain activity, primarily driven by large single-subject datasets.
+However, due to individual variation, such as anatomy, and differences in task design and scanning hardware, leveraging data across subjects and datasets remains challenging.
+In turn, the field has not benefited from the growing number of open neural data repositories to exploit large-scale deep learning.
+To address this, we develop neuroscience-informed self-supervised objectives, together with an architecture, for learning from heterogeneous brain recordings.
+Scaling to nearly 400 hours of MEG data and 900 subjects, our approach shows generalisation across participants, datasets, tasks, and even to novel subjects.
+It achieves improvements of 15-27% over state-of-the-art models and matches surgical decoding performance with non-invasive data.
+These advances unlock the potential for scaling speech decoding models beyond the current frontier.
+Machine Learning, ICML
 Figure 1: Leveraging unlabelled data using pretext tasks for speech decoding.
@@
 The tasks apply a randomly selected neuroscientifically relevant transformation $T$ to the data and the network predicts the transformation.
-We then train a linear probe on top of the pre-trained model, which remains frozen, with labelled data, achieving superior generalisation (cf.
-raw inputs) owing to the strength of the representation.
+We then train a linear probe on top of the pre-trained model, which remains frozen, with labelled data, achieving superior generalisation owing to the strength of the representation.
 1 Introduction
 In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases (Sutton, 2019).
-In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains (Jumper et al., 2021; Caron et al., 2021; OpenAI, 2023; Radford et al., 2023).
+The ability of deep learning to learn from ever-larger datasets has enabled seemingly arbitrary scaling with computation, leading to astounding advances across a diverse set of domains (Jumper et al., 2021; Caron et al., 2021; OpenAI, 2023; Radford et al., 2023).
 In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated.
-State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech.
-Yet, a number of obstacles to scale remain.
-With few exceptions at present, e.g.
-Défossez et al.
-(2023), speech decoding models tend not to train on data from more than one subject.
-Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks.
+Recent work towards brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech.
+Yet, a number of obstacles to scale remain, especially in magnetoencephalography (MEG) data.
+Current speech decoding models rarely train on multiple subjects, combine datasets, or utilise data from diverse tasks.
 Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged.
 There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example.
-But overcoming these limitations, as has begun to happen in neighbouring sub-fields, such as  Jiang et al.
-(2024), holds the promise of training models on collective, internet-scale data.
-While neuroimaging modalities such as electroencephalography (EEG) are more abundant, MEG may be a better modality for decoding as it provides a richer signal (Lopes da Silva, 2013; Hall et al., 2014).
-Given the scarcity of speech-labelled MEG data and the relative abundance of other MEG data, self-supervised learning (SSL) appears promising as it is an avenue for domains where labels are rare or hard to obtain (Balestriero et al., 2023).
-But the scale of public MEG data, while large, is still not at the volume of breakthroughs in self-supervised image and natural language processing, let alone EEG.
-Thus, SSL methods for MEG need to be highly data-efficient.
-Pretext tasks are one such method in which domain-specific self-supervised tasks are used to pre-train a model on unlabelled data by generating implicit training labels through transformations of the input in order to help a downstream task.
-We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals which we train using our pretext tasks.
-In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners.
-We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive MEG, setting the baselines for speech detection and voicing classification on this data.
-The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time.
+But overcoming these limitations holds the promise of training models on collective, internet-scale data.
+Decoding methods for MEG need to be highly data-efficient.
+While electroencephalography (EEG) data are abundant, MEG provides richer signals for decoding (Lopes da Silva, 2013; Hall et al., 2014) but are comparatively rare.
+Given the scarcity of speech-labelled MEG and the larger proportion of other MEG data, self-supervised learning (SSL) appears promising as it is an avenue for domains where labels are rare or hard to obtain (Balestriero et al., 2023).
+To data-efficiently learn from unlabelled MEG, we propose pretext training with neuroscience-informed input transformations that benefit downstream tasks.
+We use this for learning from unlabelled brain data (Figure 1) through an architecture for processing continuous multi-sensor neuroimaging signals.
+Our method provides a unified approach that enables leveraging data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners.
+We evaluate representations learned with our approach on heard speech datasets acquired with MEG, setting the baselines for speech detection and voicing classification on this data.
 Our main contributions are:
-  • A set of domain-specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
-  • A data-efficient neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
-  • A comprehensive experimental evaluation, using multiple times the volume of data in prior work, that verifies the above claims and additionally provides evidence for the existence of scaling laws when pre-training models with unlabelled MEG recordings.
+  • A domain-specific self-supervision method and a neural architecture for representation learning from MEG that unlock scaling speech decoding over multiple subjects, multiple studies, and unlabelled data;
+  • Achieving 15-27% gains over state-of-the-art self-supervised models, matching surgical self-supervised decoding non-invasively, and showing novel subject generalisation for the first time in MEG; and
+  • Demonstrating evidence for scaling laws arising from pre-training with unlabelled MEG recordings using multiple times the volume of data in prior work.
 2 Related Work
 Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments.
-This is true both in recent state-of-the-art invasive studies (Moses et al., 2021; Metzger et al., 2023; Willett et al., 2023; Chen et al., 2024) and non-invasive studies (Tang et al., 2023).
+This is true both in recent state-of-the-art invasive studies (Moses et al., 2021; Metzger et al., 2023; Willett et al., 2023; Chen et al., 2024a) and non-invasive studies (Tang et al., 2023).
 These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments.
-Focusing on semantic rather than phonetic decoding, the method developed by Tang et al.
-(2023) is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos.
-They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
-Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions.
-Csaky et al.
-(2022) find that a subset of individuals perform better when evaluated with a group-level model than with individual models.
-Exploiting audio data in a multi-modal framework, Défossez et al.
-(2023) show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated.
-However, they do not demonstrate the ability to generalise to novel subjects and must retrain their model for new datasets.
-Moreover, although they repeat the result within two MEG and two EEG datasets, Défossez et al.
-(2023) do not show any improvements for pooling data across datasets.
-Their method is also unable to incorporate data without corresponding audio labels and so they do not combine data from studies with other kinds of labels either; cf.
-Wang & Ji (2022); Duan et al.
-(2023); Wang et al.
-(2023a).
-Unfortunately, the first two of these papers included a bug in their evaluation code.
-As such, their methods may perform no better than a baseline that provides pure noise inputs to the model (Jo et al., 2024).
-[…]
-In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning (Agrawal et al., 2015; Doersch et al., 2015; Noroozi & Favaro, 2016; Larsson et al., 2016; Zhang et al., 2016; Gidaris et al., 2018).
-Until now, similar approaches have not translated to the brain decoding literature.
-However, prior work has used other methods to leverage unlabelled brain data.
-For example, Jiang et al.
-(2024) succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective.
-Although this work combined unlabelled datasets, their results studied simpler non-speech tasks with EEG.
-Wang et al.
-(2023b) used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations.
-Their impressive speech detection results were achieved with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data.
-Perhaps the closest work to ours in terms of unlocking scaling with neural data is BIOT (Yang et al., 2023).
-This is a self-supervised architecture for encoding bio-signals that is similarly capable of training with different datasets, labels, and varied numbers of sensors.
-Like the previous works, the approach tokenises signals for a transformer architecture, but instead of a masked loss it uses a contrastive pre-training objective.
-While theoretically supporting MEG, Yang et al.
-(2023) evaluate BIOT on simple ECG/EEG tasks rather than address the comparatively complex challenge of speech decoding with MEG data.
+Nevertheless, the method developed by Tang et al.
+(2023) is remarkable for showing an ability to generalise across labelled task data.
+They do not, however, use unlabelled data or show cross-subject generalisation.
+Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions (e.g.
+Anumanchipalli et al.
+(2019) and Makin et al.
+(2019) in surgical settings and Csaky et al.
+(2022) non-invasively).
+Défossez et al.
+(2023) show cross-subject generalisation for a segment identification task from participants listening to connected speech.
+However, they do not demonstrate generalisation to novel subjects and retrain their model for new datasets rather than being able to generalise across datasets or pool them.
+Their method is also unable to incorporate data without corresponding audio labels and so does not scale with other kinds of tasks.
+In general, speech decoding has centred on different kinds of speech: listening, imagining, speaking out loud, and, for paralysed patients, attempting to speak aloud.
+We focus on listening because it is easier to decode than imagined speech (e.g.
+Martin et al.
+(2014)).
+There is also evidence of functional overlap between listening and imagined speech representations in the brain (Wandelt et al., 2024), though the question of overlap has been contested (Langland-Hassan & Vicente, 2018).
+While some work on decoding text directly from heard speech tasks with MEG and EEG exist, it is unclear whether these methods perform any better than a baseline that provides pure noise inputs to the model (Jo et al., 2024).
+Non-invasive speech decoding remains a highly challenging and unsolved domain.
+Self-supervised pretext tasks have been successful in computer vision (Agrawal et al., 2015; Doersch et al., 2015; Noroozi & Favaro, 2016; Larsson et al., 2016; Zhang et al., 2016; Gidaris et al., 2018) but rarely applied to brain decoding.
+There are, however, methods that leverage unlabelled brain data in other ways (Banville et al., 2019; Kostas et al., 2021; Le & Shlizerman, 2022; Zhang et al., 2023; Yi et al., 2023; Cai et al., 2023; Ye et al., 2023; Yuan et al., 2024; Chen et al., 2024b).
+Unfortunately, most of this literature is unable to scale and harmonize heterogeneous non-invasive data.
+The most notable of these works include Wang et al.
+(2023), who learn contextualised embeddings of time-frequency input representations through masked spectrogram in-filling.
+Their impressive speech detection results were achieved with invasive neural recordings, which are comparatively rare and thus have less potential to scale than non-invasive data.
+Another, BIOT (Yang et al., 2023), learns from generic heterogeneous bio-signals with a contrastive pre-training objective, rather than masking, and applies it to ECG/EEG data.
+Notably, none of these methods optimise their objectives for speech decoding or focus on MEG.
 3 Method
-To encode continuous neuroimaging data, we introduce a neural architecture to embed heterogeneous brain signals.
-We leverage this architecture for self-supervised learning from unlabelled MEG data using a set of pretext tasks designed to generate generalisable brain representations for speech decoding.
-With this approach, we hope to replicate similar successes in computer vision (Gidaris et al., 2018; Chen et al., 2020).
+We introduce a neural architecture to embed heterogeneous brain signals.
+Then, we leverage this architecture for self-supervised learning from unlabelled MEG data using a set of pretext tasks designed to generate generalisable brain representations for speech decoding.
 3.1 Network Architecture
-Our two-stage neural network architecture (Figure 2) uses pretext tasks in pre-training to learn a representation with unlabelled brain data.
-Then, the fine-tuning stage uses this representation to learn the downstream task by training with labelled data.
+Our neural network architecture has two stages (Figure 2): pre-training with pretext tasks on unlabelled data, and training a linear probe with labelled data for downstream tasks.
 Figure 2: Architecture overview.
@@
 The network takes as input the standardised sample windows.
-To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$.
+To combine heterogeneous datasets, which have varying numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$.
 Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs (Zeghidour et al., 2022; Défossez et al., 2022).
-Specifically, our convolutional encoder adapts the implementation of the SEANet architecture (Tagliasacchi et al., 2020) used in Défossez et al.
+Specifically, our convolutional encoder adapts the SEANet architecture (Tagliasacchi et al., 2020) used in Défossez et al.
 (2022) which we describe here and as part of Figure 2.
@@
 We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$.
-Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings (where $\tau<t$), each of dimension $d_{\mathrm{backbone}}$ (i.e.
+Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings, each of dimension $d_{\mathrm{backbone}}$ (i.e.
 an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
+In the speech recognition literature, models include speaker conditioning to account for vocal and prosodic differences (Gibiansky et al., 2017).
 Just as speakers have different voices, neural responses between subjects have different characteristics.
 Consequently, individual variation leads to models that do not generalise well across subjects (Csaky et al., 2022).
-In the speech literature, models include speaker conditioning to account for these differences (Gibiansky et al., 2017).
-We take a similar approach by introducing subject conditioning.
-Zeghidour et al.
-(2022) find that conditioning is equally effective at the encoder bottleneck as in other stages of the model.
-Hence, we place ours at the cortex encoder bottleneck for simplicity.
-We use feature-wise linear modulation (FiLM) (Perez et al., 2018) as our conditioning method.
-Following the advice of Balestriero et al.
-(2023, Section 3.2), we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation.
+We address this with a similar approach to the speech literature by introducing subject conditioning using feature-wise linear modulation (FiLM) (Perez et al., 2018).
+As Zeghidour et al.
+(2022) find that conditioning is as equally effective at the encoder bottleneck as in other stages of the model, we also condition at the bottleneck.
+Following Balestriero et al.
+(2023, Section 3.2), we use a two-layer projector to alleviate misalignment between our pretext and downstream tasks in the representation.
 After the projector, linear classifiers make predictions for each of the pretext tasks.
@@
 3.2 Pretext Tasks
-Our pretext tasks are unsupervised feature learning tasks that aim to learn generalisable speech decoding features.
-Since different datasets use varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal.
+To use our architecture for pre-training, we construct pretext objectives for unsupervised learning of generalisable speech decoding features.
+These objectives are inherently agnostic to the sensor count because they operate on properties that are independent of the specific sensor arrangement.
+This key design choice enables the tasks to work seamlessly across datasets with varying numbers of sensors—a critical requirement for combining heterogeneous brain data.
+Band Hz Association
+Delta ($\delta$) .1-4 Rhythmic structure of heard speech (Luo et al., 2010)
+Theta ($\theta$) 4-8 Tracking (Luo & Poeppel, 2007) and phase-locking to the amplitude envelope of heard sentences (Peelle et al., 2012)
+Alpha ($\alpha$) 8-12 Attentional processes and the inhibition of irrelevant information (Strauß et al., 2015)
+Beta ($\beta$) 12-30 Top-down predictive coding (Bressler & Richter, 2015) which affects lexical processing (Weiss & Mueller, 2012)
+Gamma ($\gamma$) 30-70 Higher cognitive functions (e.g. memory, learning, reasoning, and planning) (Fries, 2009; Buzsáki & Wang, 2012)
+High 70+ Speech detection (Hamilton et al., 2018) and phonemic feature classification in the STG (Mesgarani et al., 2014) and the ventral sensorimotor cortex (vSMC) (Cheung et al., 2016)
+Gamma
+($\gamma^{\mathrm{high}}$)
+Table 1: Functional frequency bands in brain activity.
 Band prediction.
-In the literature, neural responses can be segmented into functional frequency bands (Giraud & Poeppel, 2012; Piai et al., 2014; Mai et al., 2016).
-Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech (Luo et al., 2010), Theta ($\theta$) waves (4–8 Hz) reliably track (Luo & Poeppel, 2007) and phase-lock to the amplitude envelope of heard sentences (Peelle et al., 2012), Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals (Strauß et al., 2015), Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding (Bressler & Richter, 2015) which affects lexical processing (Weiss & Mueller, 2012), Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g.
-memory, learning, reasoning, and planning) (Fries, 2009; Buzsáki & Wang, 2012), and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection (Hamilton et al., 2018) and phonemic feature classification in the STG (Mesgarani et al., 2014) as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) (Cheung et al., 2016).
-As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
-To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding linear predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected.
+In the literature, neural responses can be segmented into functional frequency bands (Giraud & Poeppel, 2012; Piai et al., 2014; Mai et al., 2016) (Table 1).
+Sensitivity to these frequencies would bring about functional separability in the representation space.
+Thus, to learn such representations, we train the network to classify rejected bands.
+As High Gamma is a relatively wide band we split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
+Our task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding linear predictor $f_{\mathrm{band}}$, requiring the network to classify which frequency band $\omega$ was rejected.
 This yields the loss
@@
 Phase shift prediction.
-Phase coupling between networks of neuron populations is necessary for coordinating brain activity (Fries, 2005; Vidaurre et al., 2018).
-Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature.
+Phase coupling between networks of neuron populations is necessary for coordinating brain activity (Fries, 2005; Vidaurre et al., 2018) and so phase coupling between spatially distant sensors is likely to be a useful feature.
 Supporting this insight, recent work (Jiang et al., 2024) also finds phase to be an essential component of the signal.
@@
 where $x^{A}$ is the signal scaled with $A$.
+Combined tasks.
 These pretext tasks capture complementary time- and frequency-domain properties of the signal.
 Hence, during pre-training, we combine them, creating an augmented version of the input for every pretext task by applying the matching transformation.
-We feed the augmented inputs through the network backbone and apply the corresponding classifier to predict the transformation, summing the weighted losses such that our final pre-training loss is given by
+We feed the augmented inputs through the network backbone and apply the corresponding classifier to predict the transformation, summing the weighted losses such that our final pre-training loss is
 $\mathcal{L}_{\mathrm{SSL}}=w_{1}\mathcal{L}_{\mathrm{band}}+w_{2}\mathcal{L}_{\mathrm{phase}}+w_{3}\mathcal{L}_{\mathrm{amplitude}},$ (4)
@@
 4 Experiments
-In this section, we evaluate the representations learned with our pretext tasks by measuring their ability to scale downstream performance with unlabelled data.
-This includes understanding how well they can generalise across datasets, subjects, and tasks.
+We evaluate our self-supervised representations by measuring how they scale with unlabelled data and generalise across datasets, subjects, and tasks.
 We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG (Lopes da Silva, 2013) and faster sampling rates than fMRI (Hall et al., 2014).
-We pre-train all models to completion and then fine-tune on labelled data for each task.
-In all tables and figures, we quote the receiver operating characteristic area under the curve (ROC AUC) where chance is always 0.5 regardless of class balance.
-We show the test ROC AUC at the best validation ROC AUC (early stopping) and quote uncertainty as the standard error of the mean over three seeds.
-Additionally, we state the $t$-score and $p$-value from single-sample one-sided $t$-tests against chance.
+We pre-train all models to completion and then train a linear probe on labelled data for each task.
+In all tables and figures, we quote the receiver operating characteristic area under the curve (ROC AUC) where chance is always 0.5 regardless of the class distribution.
+We show the test ROC AUC at the best validation ROC AUC (early stopping) and quote uncertainty as the standard error of the mean over up to five seeds.
 4.1 Experimental setup
 Datasets.
-Unless specified otherwise, our experiments use Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) as an unlabelled representation learning dataset for pre-training.
+In total, we use almost five times the volume of data in prior MEG work, totalling approximately 400 hours with nearly 900 subjects across pre-training and downstream training.
+Unless specified otherwise, we pre-train with Cam-CAN (Shafto et al., 2014; Taylor et al., 2017) as an unlabelled representation learning dataset.
 This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings.
-For our downstream tasks, we use two labelled heard speech MEG datasets where participants listen to short stories or audiobooks.
-Armeni et al.
-(2022) contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al.
-(2023) has 27 subjects, each recorded for 2 hours (54 hours total).
-Overall, we utilise over 200 hours of data.
-To the best of our knowledge, this is the largest volume of MEG data ever used for speech decoding.
+When aggregating datasets, we also pre-train with MOUS (Schoffelen et al., 2019), which contains 204 subjects and another 160 hours from visual and auditory tasks.
+Downstream, we use labelled heard speech MEG datasets where participants listen to short stories or audiobooks.
+We mainly focus on Armeni et al.
+(2022) which contains 3 subjects who listen to 10 hours of recordings each (30 hours total).
+We also analyse Gwilliams et al.
+(2023) which has 27 subjects, each recorded for 2 hours (54 hours total).
+The latter dataset is particularly difficult to decode from as there is very little within-subject data and it did not enforce the use of head casts to immobilise participants.
+Nevertheless, given it has many more subjects, we use this dataset to study subject generalisation.
 Preprocessing.
 Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner.
-To eliminate high-frequency muscle movement artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.5Hz to remove slow-drift artifacts.
-Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise.
+To eliminate high-frequency artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.1Hz to remove slow-drift artifacts.
+Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at 50Hz and its harmonics to account for line noise.
 Treating the low-pass filter threshold as the Nyquist frequency, we downsample the signal to twice that at 250Hz, avoiding aliasing within our band of interest.
-Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
+Finally, we detect sensor channels with significant noise and artifacts using a variance threshold and replace them by interpolating the spatially nearest sensors.
 Downstream tasks.
-We evaluate our methods with two fundamental speech decoding tasks of increasing difficulty.
-The first, speech detection, determines whether speech occurs in the auditory stimulus using the neural response.
-The second task is voicing classification.
-Given data aligned at the occurrence of a phoneme, the task is to recognise whether the phoneme is voiced or voiceless, where voicing is a binary phonetic feature that categorises whether a speech sound is associated with vocal cord vibration.
-We select these tasks as they are simpler than phoneme recognition, but are foundational because they must be solved to decode speech accurately into natural language.
+We evaluate our methods primarily on speech detection.
+This task determines whether speech occurs in the auditory stimulus using the neural response.
+This is a fundamental task in understanding speech perception and is one of the few tasks that so far show statistically significant results in highly noisy MEG signals.
+It also has direct applications to BCIs as it can be used to segment words or sentences for decoding and activate a speech BCI when a patient wishes to communicate.
+Secondarily, we also study voicing classification to demonstrate the versatility of our representations for general speech decoding tasks.
+Given data aligned at the onset of a phoneme, the task is to recognise whether the phoneme is voiced or voiceless, where voicing is a phonetic feature that categorises whether a speech sound is associated with vocal cord vibration.
+This task is also directly relevant to a speech BCI as it involves classifying phonemes which can be used to decode words.
 4.2 Learning Generalisable Representations Using Pretext Tasks
-Our first experiment investigates whether our self-supervised objectives produce generalisable representations.
-In Table 1, we show the results of pre-training models with each pretext task independently as well as together.
-Here, all of our pretext tasks lead to results that are statistically significant, and outperform a baseline fine-tuned without pre-training.
-This provides initial evidence that our tasks are helpful in speech decoding.
-Interestingly, the combination of all pretext tasks leads to better generalisation than any task on its own.
-As we hypothesised earlier, this may be because our pretext tasks capture complementary properties in time- and frequency-space, enforcing that our representation includes more salient features for speech decoding than any individual task.
-Armeni Gwilliams
-Experiment ROC AUC $t$ $p$ ROC AUC $t$ $p$
-Linear $0.559\pm 2\mathrm{e}{-4}$ $341$ $4\mathrm{e}{-6}$ $0.527\pm 7\mathrm{e}{-5}$ $379$ $3\mathrm{e}{-6}$
-BIOT + linear $0.500\pm 4\mathrm{e}{-4}$ 00 $6\mathrm{e}{-1}$ $0.499\pm 2\mathrm{e}{-4}$ $-3$ $1\mathrm{e}{+0}$
-Ours No pre-training $0.519\pm 0.002$ $8$ $7\mathrm{e}{-3}$ $0.498\pm 0.003$ 00 $7\mathrm{e}{-1}$
-+ linear Amp($\rho=0.2$) $0.602\pm 0.001$ $114$ $4\mathrm{e}{-5}$ $0.532\pm 0.005$ $6$ $1\mathrm{e}{-2}$
-Phase($\rho=0.5$) $0.603\pm 0.003$ $35$ $4\mathrm{e}{-4}$ $0.535\pm 0.003$ $12$ $3\mathrm{e}{-3}$
-Band $0.616\pm 0.003$ $44$ $3\mathrm{e}{-4}$ $0.542\pm 0.001$ $46$ $2\mathrm{e}{-4}$
-All tasks $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $\mathbf{0.543}\pm 0.003$ $13$ $3\mathrm{e}{-3}$
-Table 1: Pre-training with pretext tasks leads to better representations for decoding speech.
+Table 2 shows that our approach achieves two key feats: outperforming comparable state-of-the-art self-supervised methods by 15-27% (part C), and matching the performance of prior self-supervised methods with surgical data (11) while using only non-invasive data.
+Since non-invasive data has a much lower signal-to-noise ratio than surgical data, this is quite unprecedented.
+In the rest of this section, we analyse this table in detail.
+In part B, we show the results of pre-training models with each pretext task independently, together, and without any pre-training at all.
+Using pretext tasks (3, 4, 5) outperforms no pre-training (2).
+Interestingly, the combination of all pretext tasks (5) leads to better generalisation than any task on its own (the improvement over (3) is statistically significant).
+We conjecture that this is because our pretext tasks capture complementary properties in time- and frequency-space, ensuring that the representation includes more salient features than any individual task could encode.
+Finally, we apply Gaussian filtering to the predictions (7), smoothing out anomalies in the predicted speech envelope.
+Part / ID Model ROC AUC
+A 0 Random select .500.500.500.500
+1 Linear .$539_{\pm.002}.$
+B 2 Ours No pre-train. .$519_{\pm.002}.$
+3 + linear Amp($\rho=0.2$) .$624_{\pm.001}.$
+4 Phase($\rho=0.5$) .$615_{\pm.001}.$
+5 Band .$588_{\pm.001}.$
+6 All tasks .$630_{\pm.000}.$
+7  + smoothing .$\mathbf{.700}_{\pm.002}$
+C* 8 BIOT1 + linear .$615_{\pm.002}.$
+9 BrainBERT2 + linear .$556_{\pm.007}.$
+10 Ours (best) + linear .$\mathbf{.705}_{\pm.003}$
+11 BrainBERT2 + lin. (surgical) .$71_{\pm.06}.$
+1Yang et al. (2023) 2Wang et al. (2023)
+Table 2: Our approach outperforms all baselines and matches surgical decoding.
 In the linear-only case, we train a supervised linear classifier on the input MEG signals.
-For BIOT, we train a linear layer on top of the pre-trained BIOT representation, with the rest of the model frozen.
-Similarly, for ours, we train a linear probe on top of our pre-trained backbone with its weights frozen.
-In the no pre-training baseline, the backbone uses randomly initialised and subsequently unmodified weights.
-When all pretext tasks are used, their losses are weighted equally.
-Now, we turn to the other baselines.
-Our approach significantly outperforms the equivalent with a raw MEG input instead of a pre-trained representation (the linear experiment).
-Here, the baseline has substantially more trainable parameters because the input dimension is far larger without an encoder.
-Even with this bias favouring the experiment with the raw input, using our representation still performs better.
-We also compare our approach to BIOT (Yang et al., 2023) which is a similar state-of-the-art self-supervised method.
-When BIOT is pre-trained using exactly the same data, the fine-tuned probe fails to generalise entirely after exhaustive hyperparameter tuning.
-We put this down to three critical reasons.
-Firstly, BIOT was designed around considerably lower-dimensional signals.
-Their EEG evaluation used an order of magnitude fewer sensors than our MEG data.
-With MEG, their transformer approach requires many more channel embeddings, leading to difficulty learning the complex interactions between sensors.
-Secondly, our self-supervised objective extracts speech decoding features which is essential for solving speech decoding tasks.
-BIOT performs well on simple EEG tasks in Yang et al.
-(2023)’s evaluation, but non-invasive speech decoding is significantly more challenging.
-Together, these obstacles suggest a vast amount of data is required to learn their objective with MEG.
-Indeed, given that they pre-train with over 50 thousand hours of EEG data in their evaluation, their objective appears too general to efficiently learn a representation for speech decoding from the limited amount of MEG pre-training data (160 hours) available to us.
-This highlights the importance of data-efficiency in SSL methods for MEG.
-Among the individual pretext tasks, band prediction leads the rest.
-Perhaps this is because, by learning to discriminate between meaningful bands, the representation easily identifies phase-locking to speech onset in theta waves (Peelle et al., 2012).
-Further investigation is necessary here.
-The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search.
-We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors.
-Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018).
-In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised.
-As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.
+For ours and BrainBERT, we train a linear layer on top of a backbone pre-trained on CamCAN, with the rest of the model frozen.
+For BIOT, we use their pre-trained weights.
+In the no pre-training baseline, the backbone uses randomly initialised and frozen weights.
+In the surgical context, we quote the relevant result from (Wang et al., 2023, Table 2).
+With all pretext tasks, their losses are weighted equally.
+Next, we turn to the baselines, starting with Table 2 part A.
+Our method outperforms random selection (0) and training a linear layer with the MEG signal directly (1).
+The latter even has substantially more trainable parameters because the input dimension is larger without an encoder.
+In part C, we compare our approach to two state-of-the-art self-supervised methods.
+In each experiment, we apply Gaussian filtering as in (7).
+Although better than random, BrainBERT (9) does not generalise as well as our method (10).
+BrainBERT employs a generic masked spectrogram in-filling pre-training objective.
+While it is intended for speech decoding tasks, the pre-training objective is not designed to specifically capture features salient to neural speech processing.
+Furthermore, their method is less data-efficient because in-filling is a harder generative task compared to classification and while their methodology was developed for relatively high-fidelity intracranial recordings, the inherently lower signal-to-noise ratio of MEG presents an even greater challenge.
+Our last baseline is BIOT, which we were unable to train successfully beyond chance level using MEG data.
+Instead, in (8), we leverage publicly released weights pre-trained with over fifty thousand hours of EEG.
+However, BIOT still falls short of our pre-training method for reasons which we believe are similar to BrainBERT—its objective does not leverage neuroscientific understanding of speech processing.
+Finally, and quite remarkably, our best result matches the AUROC quoted in Wang et al.
+(2023, Table 2) who use intracranial data from heard speech (11).
+We achieved this score with non-invasive data which is typically substantially more difficult to decode due to the low signal-to-noise ratio.
+00footnotetext: Due to the large computational cost of processing embeddings for MEG data in BrainBERT, we restrict pre-training of experiments in part C to approximately 30 hours and use only subject 001 of the downstream dataset for training and evaluation.
 4.3 Scaling Speech Decoding With Unlabelled Data
@@
 We pre-train the model on increasing amounts of unlabelled data from Cam-CAN (Shafto et al., 2014; Taylor et al., 2017).
-The solid lines are the best linear fits to the data and the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
-Here, we analyse generalisation as we increase the volume of unlabelled data, analysing scaling performance on downstream tasks.
-As before, we pre-train with the combined pretext tasks.
-Figure 3 shows ROC AUC as we increase the amount of unlabelled data in pre-training up to approximately 160 hours.
-For both tasks, pre-training with any amount of data is sufficient to beat chance and there is a clear improvement in accuracy as the amount of unlabelled data increases.
+The solid lines are linear fits and the dashed lines show the volume of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
+Unlike Table 2 (7) we do not apply Gaussian filtering to the predictions for simplicity.
+The best improvements are statistically significant.
+Figure 3 shows that performance scales predictably with unlabelled data volume, following distinct patterns for different tasks and datasets.
 For speech detection on Armeni et al.
-(2022), scaling appears logarithmic in log-space; for all others, ROC AUC improves log-linearly within the data regime we study.
-In any case, adding unlabelled data has improved generalisation.
+(2022), we observe logarithmic scaling in log-space (log-log scaling), suggesting diminishing but continued returns with increased data.
+For other tasks, ROC AUC improves log-linearly, indicating robust scaling potential.
+Importantly, even our smallest pre-training dataset beats chance performance, while our largest (160 hours) continues to show gains without plateauing.
 Notably, we have scaled far beyond the data regime of prior surgical and non-surgical work and yet performance has continued to scale.
@@
 Our results also reveal several new and notable phenomena.
-Firstly, we scaled up the pre-training dataset by increasing the number of subjects.
-Since this led to consistent and almost monotonic improvements in downstream accuracy, our method is an exception to the common consensus that pooling subjects worsens generalisation.
-Secondly, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation.
-This is particularly surprising as the Armeni et al.
+We scaled up the pre-training dataset by increasing the number of subjects.
+Since this led to consistent and almost monotonic improvements in downstream accuracy, our method is an exception to the consensus that pooling subjects worsens generalisation.
+As we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation.
+This is surprising as the Armeni et al.
 (2022), Gwilliams et al.
@@
 The weaker results for Gwilliams et al.
-(2023) may be a consequence of the larger number of subjects with shorter intra-subject recordings and greater subject variation.
-These observations support the findings of other recent work such as Csaky et al.
+(2023) may be a consequence of shorter intra-subject recordings, greater subject variation, and the lack of head casts in data collection.
+These observations support the findings of work such as Csaky et al.
 (2022).
@@
 This is critical in order to widely deploy speech BCIs for new patients.
-In this experiment, we fine-tune only on Gwilliams et al.
-(2023) and hold out three subjects with which we evaluate novel subject generalisation.
+In this experiment, we use Gwilliams et al.
+(2023) as our downstream dataset because of the large number of participants, holding out three subjects with which we evaluate novel subject generalisation.
 Figure 4: Scaling unlabelled data improves novel subject generalisation.
-We fine-tune on Gwilliams et al.
+We train a linear probe on Gwilliams et al.
 (2023).
-When in-distribution, we evaluate on held-out sessions from subjects in the training set; when out-of-distribution, we evaluate on three held-out subjects.
-The solid lines are the best linear fits while the dashed lines show the amount of data used in prior surgical (Moses et al., 2021) and non-invasive (Défossez et al., 2023) work.
+When in-distribution, we evaluate on held-out sessions; when out-of-distribution, we evaluate on held-out subjects.
+The lines represent the same as in Figure 3.
+The best improvements are statistically significant.
 Figure 4 shows that scaling up the amount of unlabelled data used in pre-training not only improves accuracy on subjects previously seen, but also demonstrates a positive log-linear trend in performance for novel subjects.
 This indicates that scaling our method is an encouraging direction for resolving the challenges of subject variance faced by prior work.
-Moreover, as far as we are aware, this is the first result to demonstrate novel subject generalisation in speech decoding from MEG.
+As far as we are aware, this is the first result to demonstrate novel subject generalisation in speech decoding from MEG.
 4.5 Aggregating Unlabelled MEG Datasets
-To scale up unlabelled data further than individual studies, we must be able to combine many existing datasets.
+Given the promising scaling results with single datasets, a natural question arises: can we achieve even better performance by combining multiple MEG datasets?
+This is particularly challenging since datasets often use different scanning hardware and experimental protocols.
+Thus, it has so far not been shown in MEG.
 As a preliminary investigation, we combine two of the largest public MEG datasets: MOUS (Schoffelen et al., 2019) and Cam-CAN (Shafto et al., 2014; Taylor et al., 2017).
 In this section, we investigate how pre-training with these combined datasets affects downstream performance using the same experimental setup as Figure 3.
-Pre-training Armeni Gwilliams
-dataset Hours ROC AUC $t$ $p$ ROC AUC $t$ $p$
-Cam-CAN 159 $\mathbf{0.621}\pm 0.003$ $36$ $4\mathrm{e}{-4}$ $0.543\pm 0.003$ $13$ $3\mathrm{e}{-3}$
-MOUS 160 $0.605\pm 0.000$ $261$ $7\mathrm{e}{-6}$ $0.543\pm 0.004$ $9$ $5\mathrm{e}{-3}$
-Cam-CAN + MOUS 319 $0.611\pm 0.003$ $40$ $3\mathrm{e}{-4}$ $\mathbf{0.546}\pm 0.002$ $20$ $1\mathrm{e}{-3}$
-Table 2: Combining unlabelled datasets shows signs of outperforming single studies.
-We examine performance on the speech detection task.
-We see a small improvement when the datasets are combined on Gwilliams et al.
-(2023), but not Armeni et al.
-(2022).
+Pre-training Data Hours ROC AUC
+CamCAN1,2 $159$$630_{\pm.0001}.$
+MOUS3 $160$$614_{\pm.0004}.$
+CamCAN1,2 + MOUS3 $319$$\mathbf{.638}_{\pm.0002}$
+1Shafto et al. (2014) 2Taylor et al. (2017)
+3Schoffelen et al. (2019)
+Table 3: Aggregating unlabelled datasets outperforms single studies.
+For the first time with MEG data, we show that unlabelled pre-training data from multiple studies with different hardware profiles can be aggregated while continuing to gain the benefits of scaling.
+Combining data leads to a statistically significant $(p<0.05)$ improvement.
 The results in Table 3 show, for the first time, that combining datasets can improve performance on downstream speech decoding tasks.
-It leads to better performance on Gwilliams et al.
-(2023) compared to pre-training on either dataset alone.
-Interestingly, this was not the case for Armeni et al.
-(2022) where pre-training on Cam-CAN alone performed best.
-Combined pre-training did, however, outperform training only on MOUS.
-It is surprising that pre-training on Cam-CAN was better than pre-training on MOUS when evaluating on Armeni et al.
-(2022) given that MOUS and Armeni et al.
-(2022) both used speech tasks and were acquired on the same MEG scanner.
+It leads to better performance compared to pre-training on either dataset alone.
+It is surprising that pre-training on Cam-CAN was better than pre-training on MOUS given that MOUS and the downstream dataset both used speech tasks and were acquired on the same MEG scanner.
 Cam-CAN, by contrast, did not use a speech task and was acquired on a different MEG scanner.
-We hypothesise that the better results for Cam-CAN are due to it being a cleaner dataset.
-During our experiments, we found that data quality, even among unlabelled data, can have a significant affect as artefacts in recordings disrupt learning.
+We hypothesise that the better results for Cam-CAN are due to it being cleaner.
+During our experiments, we found that data quality, even among unlabelled data, can have a significant effect as artefacts in recordings disrupt learning.
 While the combination of the two datasets includes far more hours of data than any prior work on deep learning with MEG, further work needs to be done to aggregate more datasets.
-Here, we were limited by compute budget.
-Increasing the number of datasets could enable the network to eventually always improve over the best singular dataset.
+Here, we were limited by compute budget and data availability.
+Increasing the number of datasets (e.g., by including EEG too) could further improve results.
 Just as increasing the number of subjects (rather than only within-subject data) improves novel subject generalisation, a larger number of datasets may be key to scaling results when datasets are aggregated in pre-training.
@@
 Although our results are significant in demonstrating a viable path forward to scale up speech BCIs, there remain a number of limitations to the present work.
-We focused here on two downstream tasks: speech detection and voice classification.
+We focused here on two downstream tasks: speech detection and voicing classification.
 Ultimately, we would like to expand this work to predict full transcripts from brain recordings (i.e.
@@
 Speech detection has played an important role in the development of full brain-to-text in a surgical context (Moses et al., 2021) and we hope may play a similar role for non-invasive methods.
-Prior work has further used voice classification as a stand in for phoneme classification (Gwilliams et al., 2022), and we have been able to improve on these results here.
-In future work, we would like to expand this to all English phonemes.
-Secondly, while we have been able to demonstrate the utility of a few pretext tasks, we do not claim to have exhausted the full set of useful tasks.
+In future work, we would also like to expand classification to all English phonemes as a step towards full transcript decoding.
+Additionally, while we have been able to demonstrate the utility of a few pretext tasks, we do not claim to have exhausted the full set of useful tasks.
 Rather, we conjecture that more useful pretext tasks remain to be found and believe a useful avenue of research will be into other input representations for brain recordings.
-For example, this paper did not make use of spatial features.
+For example, this paper did not make use of spatial features when the geometry of a scanner’s sensor configuration is strongly correlated with the area of the brain from which the signal is derived.
 Another limitation is our emphasis on heard speech over other types of speech, such as attempted or imagined speech.
 We hypothesise that the same methods presented here will generalise to these other varieties of speech, though this has yet to be shown.
-But, perhaps the biggest limitation of the present work is that, while it surpasses the amount of data used in other studies, it remains to be seen how much speech decoding tasks can be improved by scaling up the number of datasets used in training.
+Perhaps the biggest limitation of the present work is that, while it surpasses the amount of data used in other studies, it remains to be seen how much speech decoding tasks can be improved by scaling up the number of datasets used in training.
 In sharing this work now, we believe that the current proof of concept will be sufficiently impactful to the field as we continue to actively scale up the datasets that we can leverage.
 5 Conclusion
-Ultimately, solving speech decoding could transform the lives of patients with severe communication difficulties.
-This promise has not yet materialised because the field has been blocked by its inability to scale up data to leverage deep learning.
+Speech decoding from the brain has been limited by the field’s inability to scale up data to leverage deep learning.
 Prior methods have been unable to aggregate data across different datasets, labels, or subjects to scale up because of heterogeneity in recording hardware, experiment design, and participants.
@@
 But until now, no one has developed a general solution.
-We provided a unified method that leverages unlabelled recordings data-efficiently using generic pretext tasks that shows that all of these problems can be solved.
-We verified this with experiments showing that our method not only scales with heterogeneous data but even generalises across datasets, subjects, and tasks.
+We present a unified solution through data-efficient, self-supervised pretext tasks that overcome these fundamental scaling challenges.
+Our experiments demonstrate not just scaling with heterogeneous data, but generalisation across datasets, subjects, and tasks.
+They also show significant improvements of up to 27% compared to the prior state-of-the-art and even provide evidence of matching surgical decoding performance.
 Our method unlocks the potential of the bitter lesson, providing a general method to exploit more computation by using more data.
@@
 If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
-Ethics Statement
-In this work, we use data from studies that involve human subjects (Armeni et al., 2022; Gwilliams et al., 2023; Shafto et al., 2014; Taylor et al., 2017; Schoffelen et al., 2019).
-These datasets are public, cited, and have their own ethical approvals.
-The documentation for these is available with the publications for the respective datasets.
-While there are clear positive impacts, we acknowledge that insights from neural speech decoding research may not all be beneficial.
-Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g.
-stammering).
-As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants.
-Nevertheless, the maturity of this technology could also present potential negative societal impacts.
-For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals.
-Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent.
-Moreover, there are possible long horizon effects associated with speech decoding research.
-Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society.
-In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality.
-Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse.
-Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.
-Reproducibility Statement
-In the supplementary materials, we have provided an anonymised code repository with instructions for reproducing our main experiments.
-We also include details on experiment design and setup (Section 4.1 and Appendix A), hyperparameters (Appendix B), and compute (Appendix C).
-While we attempt to be exhaustive with these details, any information not found directly in the main body or appendices can be located in the supplementary materials.
+Impact Statement
+This work uses publicly available datasets from human studies (Armeni et al., 2022; Gwilliams et al., 2023; Shafto et al., 2014; Taylor et al., 2017; Schoffelen et al., 2019), each with their own ethical approvals and documentation available in their respective publications.
+Neural speech decoding research has transformative potential for healthcare and assistive technology.
+Advances could help paralysed patients communicate freely and assist those with communication difficulties.
+By developing non-invasive methods, the field opens up the possibility of broader access to these technologies without the risks of surgical implants.
+However, we acknowledge potential societal risks as this technology matures:
+  • Privacy and Data Protection: Brain signals contain highly sensitive personal information, raising concerns about data security and individual privacy.
+  • Consent and Misuse: Advanced decoding capabilities could enable unauthorized access to neural information, requiring robust safeguards against exploitation.
+  • Societal Impact: Widespread adoption could affect privacy norms around inner speech, while unequal access could exacerbate existing inequalities.
+We focus specifically on decoding heard speech rather than inner speech, limiting potential misuse.
+Nevertheless, we recognize that advances in heard speech decoding contribute to the broader development of neural decoding technology.
+We encourage the research community to actively engage with these ethical considerations as the field progresses.
 Appendix A Experiment Details
+Pre-training data.
 We pre-train with non-overlapping sample windows from all subjects and sessions.
@@
 Each seed uses a different set of subjects to reduce negative effects from outlier subjects.
-When fine-tuning with Armeni et al.
-(2022), we hold out session 010 from all subjects during training and validation, using this for evaluation.
+Labelled training data.
+When training with Armeni et al.
+(2022), we hold out session 009 for validation and 010 for testing.
 Similarly, when fine-tuning with Gwilliams et al.
-(2023), we hold out session 1 from subjects 23, 24, 25, 26, and 27, using these sessions for evaluation only.
+(2023), we hold out task 1 from subjects 23, 24, 25, 26, and 27, using these sessions for evaluation only.
 As there is limited within-subject data in the latter dataset, we did not hold out a session from all subjects as before.
@@
 Both sessions for each task are repeats of the task.
-This means that while the recording itself is unseen, in this dataset, it is possible that heldout sessions use tasks that may have been seen in the training set.
-In all experiments, we always fine-tune to completion (usually around 300 epochs), taking the test metric at the best validation loss (early stopping).
-We use three randomly selected seeds for each pre-training and corresponding fine-tuning run.
-For speech detection, since our encoder reduces the temporal dimension from 125 samples (the number of samples in a 0.5 second window with a sample rate of 250Hz) down to 5 embeddings, we downsample our speech detection labels to match using PyTorch’s torch.nn.functional.interpolate.
-Therefore, each speech detection label represents a 0.1 second period of time.
+This means that while the recording itself is unseen, in this dataset, it is possible that held-out sessions use stimuli that may be shared.
+BrainBERT adaptation.
+As BrainBERT is designed for single-electrode representations, for a fair comparison, to take into account all sensors, we ensured our linear classifier is applied to a concatenated embedding over all sensors.
+This is a large vector and leads to very computationally expensive training and is the reason we had to reduce the pre-training and downstream data in part C of Table 2.
+Statistical testing. Significance was determined using one-sided $t$-tests with $p<0.05$ as the threshold for significance.
 Appendix B Hyperparameters
@@
 Table 4: Experimental hyperparameters.
+Why $\rho$?
+The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search.
+We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors.
+Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018).
+In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised.
+As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.
 Appendix C Compute Resources

### v5

@@
 But overcoming these limitations holds the promise of training models on collective, internet-scale data.
+https://pnpl.robots.ox.ac.uk/bbl
 Decoding methods for MEG need to be highly data-efficient.
@@
 Supporting this insight, recent work (Jiang et al., 2024) also finds phase to be an essential component of the signal.
-To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho$ of the sensors.
+To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho\in[0,0.5]$ of the sensors.
 Applying this shift to random sensors is critical since sensors are placed in different positions, capturing different regions of the brain.
@@
 Thus, this pretext task focuses on learning representations that encode relative sensor amplitude differences.
-Similar to the phase shift task, we select a random proportion of the sensors $\rho$ and apply a discrete random amplitude scaling coefficient $A\in[-2,2]$, discretised into 16 scaling factors, to the signal.
+Similar to the phase shift task, we select a random proportion of the sensors $\rho\in[0,0.5]$ and apply a discrete random amplitude scaling coefficient $A\in[-2,2]$, discretised into 16 scaling factors, to the signal.
 The objective is to predict the scaling factor, leading to the loss
@@
 $\mathcal{L}_{\mathrm{SSL}}=w_{1}\mathcal{L}_{\mathrm{band}}+w_{2}\mathcal{L}_{\mathrm{phase}}+w_{3}\mathcal{L}_{\mathrm{amplitude}},$ (4)
-where $w_{i}$ is a constant coefficient for each self-supervised loss.
+where $w_{i}$ is a constant coefficient for each loss.
 4 Experiments
@@
 9 BrainBERT2 + linear .$556_{\pm.007}.$
+10 EEGPT3 + linear .$602_{\pm.006}.$
 11 Ours (best) + linear .$\mathbf{.705}_{\pm.003}$
 12 BrainBERT2 + lin. (surgical) .$71_{\pm.06}.$
-1Yang et al. (2023) 2Wang et al. (2023)
-Table 2: Our approach outperforms all baselines and matches surgical decoding.
-In the linear-only case, we train a supervised linear classifier on the input MEG signals.
+1Yang et al. (2023) 2Wang et al. (2023) 3Wang et al. (2024)
+Table 2: Our approach surpasses baselines in speech detection by up to 27% and matches surgical decoding.
+For linear, we train a supervised linear classifier on the MEG signals.
 For ours and BrainBERT, we train a linear layer on top of a backbone pre-trained on CamCAN, with the rest of the model frozen.
@@
 In the no pre-training baseline, the backbone uses randomly initialised and frozen weights.
-In the surgical context, we quote the relevant result from (Wang et al., 2023, Table 2).
-With all pretext tasks, their losses are weighted equally.
+In the surgical context, we quote the result from Wang et al.
+(2023, Table 2).
+With all pretext tasks, losses are weighted equally.
 Next, we turn to the baselines, starting with Table 2 part A.
@@
 Furthermore, their method is less data-efficient because in-filling is a harder generative task compared to classification and while their methodology was developed for relatively high-fidelity intracranial recordings, the inherently lower signal-to-noise ratio of MEG presents an even greater challenge.
-Our last baseline is BIOT, which we were unable to train successfully beyond chance level using MEG data.
-Instead, in (8), we leverage publicly released weights pre-trained with over fifty thousand hours of EEG.
-However, BIOT still falls short of our pre-training method for reasons which we believe are similar to BrainBERT—its objective does not leverage neuroscientific understanding of speech processing.
+Our last baselines are BIOT (8) and EEGPT (10).
+We leverage publicly released weights pre-trained with thousands of hours of EEG.
+However, both still fall short of our pre-training method for reasons which we believe are similar to BrainBERT—their objectives do not leverage neuroscientific understanding of speech processing.
 Finally, and quite remarkably, our best result matches the AUROC quoted in Wang et al.
@@
 3Schoffelen et al. (2019)
-Table 3: Aggregating unlabelled datasets outperforms single studies.
-For the first time with MEG data, we show that unlabelled pre-training data from multiple studies with different hardware profiles can be aggregated while continuing to gain the benefits of scaling.
-Combining data leads to a statistically significant $(p<0.05)$ improvement.
+Table 3: Aggregating unlabelled datasets outperforms single studies in speech detection.
+For the first time with MEG, we show that unlabelled pre-training data from multiple studies with different hardware profiles can be aggregated while gaining the benefits of scaling.
+Combining data leads to a significant $(p<0.05)$ improvement.
 The results in Table 3 show, for the first time, that combining datasets can improve performance on downstream speech decoding tasks.
@@
 If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
+Acknowledgements
+We would like to thank Botos Csaba for many early insightful discussions which helped shaped the direction of this work.
+In alphabetical order, thanks also to Mats W.J.
+van Es for technical assistance with the OSL library, Yonatan Gideoni for advice on data splits, Minqi Jiang for an encouraging conversation on scaling unsupervised representation learning, Brian Liu for technical contributions which did not reach the final paper, and Miran Özdogan for reviewing a draft of this work.
+Finally, we thank the rest of the PNPL group for their continued and unwavering support.
+The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558.
+[…]
 Impact Statement
@@
 This means that while the recording itself is unseen, in this dataset, it is possible that held-out sessions use stimuli that may be shared.
-BrainBERT adaptation.
+Adapting models for more sensors.
 As BrainBERT is designed for single-electrode representations, for a fair comparison, to take into account all sensors, we ensured our linear classifier is applied to a concatenated embedding over all sensors.
 This is a large vector and leads to very computationally expensive training and is the reason we had to reduce the pre-training and downstream data in part C of Table 2.
+We similarly concatenate sets of 19 sensors when evaluating EEGPT.
 Statistical testing. Significance was determined using one-sided $t$-tests with $p<0.05$ as the threshold for significance.
@@
 Table 4: Experimental hyperparameters.
-Why $\rho$?
-The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search.
-We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors.
-Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018).
-In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised.
-As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.
+[…]
 Appendix C Compute Resources